### v1

# Family Matters: Language Transfer and Merging for Adapting Small LLMs to Faroese

Abstract
We investigate how to adapt small, efficient LLMs to Faroese, a low-resource North Germanic language. Starting from English models, we continue pre-training on related Scandinavian languages, either individually or combined via merging, before fine-tuning on Faroese. We compare full fine-tuning with parameter-efficient tuning using LoRA, evaluating their impact on both linguistic accuracy and text comprehension. Due to the lack of existing Faroese evaluation data, we construct two new minimal-pair benchmarks from adapted and newly collected datasets and complement them with human evaluations by Faroese linguists. Our results demonstrate that transfer from related languages is crucial, though the optimal source language depends on the task: Icelandic enhances linguistic accuracy, whereas Danish boosts comprehension. Similarly, the choice between full fine-tuning and LoRA is task-dependent: LoRA improves linguistic acceptability and slightly increases human evaluation scores on the base model, while full fine-tuning yields stronger comprehension performance and better preserves model capabilities during downstream fine-tuning.
1 Introduction
While large language models (LLMs) excel in English and other high-resource languages, low-resource languages lag behind: model quality is tightly linked to data availability Robinson et al. (2023); Li et al. (2024), and even basic comprehension may fail Court and Elsner (2024). Coverage of these languages typically requires the largest available models, if it works at all.
In this paper, we focus on Faroese, a North Germanic language spoken by around 70,000 people, mostly in the Faroe Islands. Training data is scarce, with only 95 million words available in the deduplicated Fineweb-2 dataset Penedo et al. (2024). Faroese is particularly interesting for studying transfer from related languages. In the Nordic context, it has been described as the central Nordic language Torp (1998), reflecting its typological position in relation to the other Scandinavian languages. Derived from Old Norse, shaped by historical development and sociopolitical circumstances, modern Faroese shares characteristics with all the other Scandinavian languages.
We investigate methods for adapting small generative models to Faroese through continued pre-training, with particular emphasis on leveraging transfer from related languages. Prior work has demonstrated that syntactic similarity is a strong predictor of transfer success Chang et al. (2024), and for Faroese, encoder models benefit notably from transfer from Icelandic and other Scandinavian languages Snæbjarnarson et al. (2023). Beyond direct transfer, we explore parameter merging (Wortsman et al., 2022; Ilharco et al., 2023; Yadav et al., 2023), which enables the combination of models fine-tuned on different languages. This approach allows us to control the relative influence of each source language, a flexibility that would be more costly to achieve via multilingual training. Our goal is to balance the close linguistic similarity of Icelandic with the larger data resources available for mainland Scandinavian languages. To this end, we merge models trained on various Scandinavian languages and subsequently fine-tune the merged models on Faroese. We also compare two fine-tuning strategies for the continued pre-training: full fine-tuning and parameter-efficient fine-tuning with LoRA (Hu et al., 2022). Prior evidence suggests that full fine-tuning provides higher accuracy, while LoRA better preserves previously learned skills and diversity in text generation (Biderman et al., 2024). Based on this, we hypothesize that full fine-tuning will yield stronger linguistic performance in Faroese, but with a higher risk of losing general reasoning and knowledge abilities.
A major challenge for low-resource languages is the scarcity of evaluation data. To address this, we take two steps. First, we introduce two minimal-pair evaluation suites for small LLMs, combining pre-existing, newly collected, and adapted datasets: FoBLiMP, which targets syntactic and linguistic acceptability, and FoBCoMP, which targets text comprehension. Second, we perform expert human evaluations (carried out by Faroese linguists) on two tasks, one with and one without low-resource downstream task fine-tuning.
Our results indicate that transfer from related languages is crucial. Transfer from Icelandic yields the best performance in the linguistic probes in FoBLiMP, whereas higher-resource mainland Scandinavian languages contributes more to the text comprehension probes in FoBCoMP. Language merging shows promise for combining these advantages, but improvements are not consistent. Contrary to our expectations, LoRA outperforms full fine-tuning on linguistic acceptability in both automatic and human evaluations. In contrast, full fine-tuning achieves better results on comprehension tasks and coupled with downstream fine-tuning.
Our research questions are the following:
  1. RQ1
    What is the effect of transfer from (to various degrees) related languages?
  2. RQ2
    Does merging various related languages have a benefit over choosing the closest neighbor?
  3. RQ3
    What are the differences between fine-tuning all parameters of the base model and performing adaptation with LoRA?
2 Related Work
Language Adaptation
Post-hoc adaptation transfers knowledge from a high-resource or multilingual model to a new language, often using parameter-efficient methods. Early work focused on encoder models (Pfeiffer et al., 2020; Ansell et al., 2022; Ebrahimi and Kann, 2021). Yong et al. (2023) found that full fine-tuning works best for smaller models (e.g., 560M), while adapters are better for larger ones (up to 7.1B), but despite using a generative model, they do not evaluate on generative tasks. Razumovskaia et al. (2024) show that continued pre-training with LoRA (similar to our setup) improves linguistic quality of generations, though not few-shot ability. Instruction tuning on translated data is another option for language adaptation (Muennighoff et al., 2023b), but it is less reliable for small languages, and even correct translations can produce “translationese” (Gellerstam, 1986).
Selecting effective transfer languages is crucial for language adaptation. Chronopoulou et al. (2023) show that sharing adapter parameters across related languages improves machine translation for low-resource languages. Similarly, Faisal and Anastasopoulos (2022) find that linguistically informed adapter designs benefit unseen languages. In machine translation, transfer success often depends more on source corpus size and subword overlap than on broader linguistic similarity (Lin et al., 2019). Pairing a low-resource language with a typologically similar, higher-resource language can outperform using all available languages (Neubig and Hu, 2018). However, they focus on translation from a low-resource language; a setup that does not require generation in the small language.
Faroese Typology and Its Implications
Faroese is classified as Insular Scandinavian with Icelandic, while Norwegian, Swedish, and Danish are Mainland Scandinavian. This division reflects differences in lexicon, morphology, and syntax, though Faroese also shares traits with Mainland Scandinavian languages Thráinsson et al. (2012). Lexically, Faroese shares words and cognates with Icelandic and Danish, but no quantitative analysis exists Jacobsen (2021, 2022). Morphologically, it is more similar to Icelandic, with some overlap with Norwegian Torp (1998). Syntactically, Faroese is closer to Mainland Scandinavian languages Ussery and Petersen (2023); Debess (2017); Petersen (2010); Petersen and Heycock (2017); Sandøy (2005). Because of this mixed profile, identifying a single closest neighbour language is difficult. In our experiments, we fine-tune using data from all Scandinavian languages, with particular attention to Icelandic, to capture the different typological influences.
Evaluation for low-resource languages and small models
is challenging because there is little task-specific data. On the other hand, few-shot evaluations often focus on knowledge-intensive tasks or tasks that require structured outputs, making them out of scope for small models. Perplexity is the simplest measure of a model’s fit to a language, and only requires held-out text. In low-resource languages, however, fine-grained tokenization from common tokenizers reduces its usefulness (Oh and Schuler, 2024). An alternative is information parity, which compares the negative log-likelihood of a target-language text to its English version (Tsvetkov and Kipnis, 2024). This metric correlates with downstream performance but requires parallel data. Another option is translated benchmarks, but these miss culture-specific content (Chen et al., 2024), introduce translation errors, and produce “translationese”, making the target language artificially easier for the model.
Minimal pairs consist of two sentences that differ slightly, with one correct and one incorrect. Models are expected to assign higher probability to the correct sentence, capturing implicit preferences (Marvin and Linzen, 2018; He et al., 2025). Classical examples include subject-verb agreement (Linzen et al. (2016): “The key is on the table.” vs. “The key are on the table.”) and negative polarity items (Marvin and Linzen (2018): “No students have ever lived here.” vs. “Most students have ever lived here.”). BLiMP (Warstadt et al., 2020) provides a wide range of syntactic minimal pairs, with multilingual extensions like MultiBLiMP (Jumelet et al., 2025) including Faroese. COMPS (Misra et al., 2023) tests semantic knowledge (e.g., “A robin can fly.” vs. “A penguin/table can fly.”), and its multilingual extension (He et al., 2025) exists, though it does not cover Faroese.
Recent work for Faroese has focused on machine translation Scalvini and Debess (2024); Scalvini et al. (2025a); Simonsen and Einarsson (2024); Debess et al. (2025), combining automatic and human assessments. Automatic metrics such as BLEU or chrF prove shallow and fail to capture Faroese-specific nuances Scalvini et al. (2025b). Embedding-based metrics, such as BERTScore via FoBERT Snæbjarnarson et al. (2023), are emerging but limited and rarely validated against human judgment. Human evaluation remains essential. Beyond MT, small Faroese evaluation datasets exist for sentiment analysis Debess et al. (2024) and question answering Simonsen et al. (2025).
3 Experimental Setup
In this section, we introduce our experimental setup including model training (Section 3.1), merging (Section 3.2), collection of data for the automatic evaluation (Section 3.3), and the human evaluation process and data (Section 3.4). [^1] We make models and data available in a huggingface collection: huggingface.co/collections/jekunz/faroese-adaptation-68d6aea566b16d4f57180682.
3.1 Training
We use the two smaller SmolLM2 (Allal et al., 2025) models (135M and 360M parameters) as they are fully open, including their training data, and well-trained for their size on an English corpus. Although their size limits performance on knowledge-intensive tasks, it allows us to continually pre-train and compare them in different setups across substantial corpora. We experiment with two adaptation setups to answer RQ3: full-parameter fine-tuning and LoRA fine-tuning. We train for 5 epochs on the Faroese corpus (following Muennighoff et al. (2023a)’s scaling law for data-constraint training); we do not repeat data for other languages. Training details can be found in Table 6 in Appendix A. We train on the deduplicated Fineweb-2 (Penedo et al., 2024) portions for the Scandinavian languages, containing 27B tokens for Danish, 25B for Swedish, 30B for Norwegian-Bokmål, 1.6B for Icelandic, 495M for Norwegian-Nynorsk, and 95M for Faroese. Due to resource constraints, we limit the corpora for Swedish, Danish and Norwegian (Bokmål) to 4B tokens. We perform sequential continued pre-training for RQ1, first on an individual transfer language, then on the Faroese. We do not merge the data of the source languages with the Faroese data because this may introduce language mixing issues (Li et al., 2025).
3.2 Merging
Merging provides an efficient alternative to multi-task training by combining different fine-tuned models into one checkpoint, often improving generalization (Wortsman et al., 2022; Yadav et al., 2024). Task Arithmetic merging (Ilharco et al., 2023) computes task vectors as the difference between a fine-tuned model and its base, averages them, and adds the result back to the base model. TIES (Yadav et al., 2023) refines this by keeping only the most influential parameters, resolving sign conflicts, and using a disjoint mean to reduce interference. We apply TIES to languages instead of tasks, using Mergekit (Goddard et al., 2024): starting from SmolLM, we fine-tune on each language to obtain language vectors, merge them, and train the resulting model on Faroese. Since we continue pre-training on Faroese after merging and have limited resources, we select three promising merges that cover different language mixes and weightings: Mergeeq where we merge all five models equally (with pre-normalization weight 1 and density 0.5), Mergeis+ where we merge with bias towards Icelandic (weight 1 for Icelandic, 0.5 for all others, density 0.5), and Merge2, where we merge only two models: Icelandic and Danish (both with weight 1 and density 0.5), as Danish is the mainland Scandinavian language with the lowest perplexity after Faroese continued pre-training (see Table 3(b)).
3.3 Automatic Evaluation
We evaluate the perplexity on the validation set of the Faroese portion of Fineweb-2 both zero-shot and after continued pre-training on Faroese. In addition, we introduce two benchmarks: FoBLiMP for linguistic acceptability probes, and FoBCoMP for text comprehension probes. We report results on the original SmolLM models as a baseline.
FoBLiMP
To probe zero-shot linguistic skills, we use minimal pairs with one correct and one corrupted sentence, measuring the percentage of times the model assigns higher probability to the correct sentence. This collection is called FoBLiMP (Faroese Benchmark of Linguistic Minimal Pairs).
To evaluate subject-verb agreement, we use the Faroese portion of MultiBLiMP (Jumelet et al., 2025), containing 232 sentences. ScaLA (Nielsen, 2023) contains sentences corrupted by swapping or deleting words. Originally a binary classification task, we convert it to minimal pairs by realigning correct and incorrect sentences using Levenshtein distance ($\geq 0.85$), with unmatched samples added manually. Concatenating all subsets gives 552 pairs for flip_neighbours and 601 pairs for delete. GermDetect Michael and Horbach (2025) provide automatically corrupted sentences with verb placement errors. After removing pairs with no corruption, we obtain 2,026 pairs. As Faroese allows flexible word order, some corruptions are grammatical, but we conclude from an inspection that the original sentence is mostly more common. We also construct minimal pairs from a human evaluation in Scalvini et al. (2025a), where two raters annotated errors in English-to-Faroese translations from four models. We pair translations with an error difference of at least 2, keeping those with no more than four errors in the better translation and excluding translations containing foreign scripts. This yields 680 pairs.
FoBCoMP
Evaluating small LLMs in text comprehension is particularly challenging because evaluations often mix formal competence (e.g., grammar) with functional competence (e.g., following prompts) (Kydlíček et al., 2024). Limited fine-tuning data further complicates comparisons. To address this, we also use text comprehension probes in a minimal-pair format. We introduce a set of five probes, called FoBCoMP (Faroese Benchmark of Text Comprehension Minimal Pairs).
We adapt the Faroese news sentiment dataset (Debess et al., 2024) (original labels: positive, negative, neutral) into minimal pairs by adding a sentiment-bearing sentence (“Hetta er gott/ringt”). Neutral labels are excluded as initial experiments showed that words such as “neutral” are never the most probable choice. We evaluate sentence- and article-level samples, keeping only items annotators of the original dataset agreed on, resulting in 91 sentence-level (55 positive, 36 negative) and 84 article-level (51 positive, 33 negative) pairs. Using the same dataset, we filter GPT-4-assigned topic labels confirmed by a human. Minimal pairs consist of one correct topic and one incorrect topic (not assigned to the article), with related-topic pairs curated to make the task realistic (e.g., Local News vs. International News). This yields 234 topic classification pairs. We also adapt the extractive QA dataset FoQA (Simonsen et al., 2025) (2,000 Faroese question–context–answer triplets) into minimal pairs via two methods: (1) Dataset Shuffling: Replacing the correct answer passage with an incorrect but plausible passage from another sample within the context, matching token length. This creates 21,867 pairs. (2) GPT-4 Adversarial Answers: Generating one alternative incorrect answer per sample that is also a span in the dataset, matching token length when possible. Exact length matches occurred for 611 answers; deviations were an average of 1.69 tokens longer than the correct answers. This yields 2,000 pairs.
3.4 Human Evaluation
As we do not have evaluation sets to assess fine-grained properties for Faroese generation, we perform a human evaluation to assess output quality. The latter requires subjective judgment of linguistic quality, naturalness, and contextual appropriateness. As human evaluation is time-intensive, we focus on four 360M models, which perform among the best. These models vary along two dimensions: full tuning versus LoRA, and transfer from Icelandic versus Mergeis+. The evaluators are two native Faroese speakers, both trained linguists with extensive experience in NLP model evaluation.
Sentence Continuation
Given that we are working with small base models, we chose a simple generation task: sentence continuation. Models are prompted with sentences from a small manually compiled corpus derived from academic papers and local news (not in FineWeb2), with the last words removed and a trailing space. All models produce running text, enabling comparison of output quality. We evaluate outputs for linguistic quality across four subdimensions, scored 0–5: Lexical correctness (valid Faroese words and avoidance of lexical hallucinations), grammatical accuracy (morphological and syntactic correctness, including spelling and typography), Semantic coherence (meaningful, logically consistent content) and fluency/naturalness (native-like expression). The annotators each evaluated 400 continuations (100 prompts, 4 models) We apply a token cut-off of 100 and no penalty for incomplete last sentences.
Summarization
As zero-shot summarization did not result in reasonable summaries, we create a synthetic dataset for fine-tuning and manually evaluate the summaries generated by the models. The synthetic dataset is compiled with 150 authentic texts in the same domains as the evaluation set (academic, news, blog) paired with summaries generated by Claude Sonnet 4222For fine-tuning details, see Table 7 in Appendix A.. For evaluation, we selected 50 source texts (not part of training text) and generated summaries from all four models, resulting in 200 summary-source pairs. Two annotators conducted a blind evaluation using two criteria: task completion (to what extent did the model solve the task) and linguistic quality, each on a 0–5 scale. We apply a token cut-off of 400 and no penalty for incomplete final sentences.
4 Results and Discussion
4.1 Benchmarks
We first give an overview of the benchmarks results with respect to their difficulty and reliability. FoBLiMP results are shown in Table 1. Models perform well on most linguistic acceptability probes, suggesting these tasks are relatively easy. The main exception is Translation Pairs, where scores are lower, likely due to noise as translation error counts do not always reflect linguistic quality: an error can reflect unrelated aspects such as incorrect content compared to the source sentence.
Full LoRA 135M 360M 135M 360M En 95.25 96.55 96.12 96.55 +Da 96.98 95.25 98.27 96.12 +Is 97.41 99.13 97.41 98.27 +NoB 96.12 100.0 96.98 96.98 +NoN 96.12 98.27 95.68 97.41 +Sv 96.12 98.27 96.98 96.98 Mergeeq 97.41 98.27 96.55 95.25 Mergeis+ 97.84 97.84 97.41 96.55 Merge2 97.84 97.41 97.41 99.13
(a) Subject-verb agreement (MultiBLiMP).
Baseline: 66.81 (135M), 70.68 (360M)
Full LoRA 135M 360M 135M 360M 93.65 94.56 92.75 93.84 94.74 94.02 95.10 96.01 93.84 95.65 94.38 96.01 94.38 95.10 95.28 95.10 93.65 94.92 94.02 94.20 95.28 94.20 95.65 94.92 94.56 94.38 96.73 95.47 94.20 95.28 95.10 95.10 95.28 95.10 93.84 95.28
(b) ScaLA: flip_neighbors.
BL: 59.96 (135M), 62.50 (360M)
Full LoRA 135M 360M 135M 360M 87.18 93.17 84.52 88.51 90.84 93.01 90.18 92.84 89.51 94.00 90.18 95.34 90.34 92.34 90.68 94.50 88.51 93.17 89.01 91.01 89.85 93.34 90.51 93.17 90.18 94.34 89.35 93.01 92.01 94.00 89.01 93.17 90.68 93.51 90.34 91.84
(c) ScaLA: delete.
BL: 65.39 (135M), 67.38 (360M)
Full LoRA 135M 360M 135M 360M En 93.48 95.36 92.54 94.66 +Da 94.27 95.11 94.91 95.75 +Is 94.27 95.16 94.57 95.75 +NoB 95.06 95.31 94.91 95.55 +NoN 93.97 94.91 94.47 95.01 +Sv 94.61 95.85 95.36 95.80 Mergeeq 94.66 95.80 95.16 95.31 Mergeis+ 94.57 95.60 95.06 95.26 Merge2 94.52 95.85 94.37 95.60
(d) Verb placement (GermDetect).
Baseline: 52.22 (135M), 57.94 (360M)
Full LoRA 135M 360M 135M 360M 70.44 75.00 68.97 74.70 75.58 78.23 75.88 76.76 73.23 77.35 75.44 78.67 75.14 76.17 75.44 77.79 70.73 75.44 71.17 74.70 76.32 76.17 75.88 74.11 76.02 75.88 75.00 75.44 76.91 77.20 74.55 73.97 74.55 77.20 75.44 75.44
(e) Translation Pairs.
BL: 42.64 (135M), 46.32 (360M)
Full LoRA 135M 360M 135M 360M 88.00 90.92 86.78 89.65 90.48 91.12 90.86 91.49 89.65 92.25 90.39 92.80 90.20 91.78 90.65 91.98 88.59 91.34 88.87 90.46 90.43 91.55 90.87 90.99 90.56 91.73 90.55 90.89 91.10 91.98 90.22 90.81 90.57 91.81 90.28 91.45
(f) Mean of the individual FoBLiMP scores in 1(a)-1(e).
Table 1: Linguistic probes on datasets included in FoBLiMP: Percentage of samples where a higher probability was assigned to the original than to the corrupted sample.
FoBCoMP results are shown in Table 2. Scores are lower and more mixed than for FoBLiMP, reflecting the tasks’ difficulty for small models. In sentiment analysis, 135M LoRA models perform poorly, with little improvement over the base model, and full fine-tuning shows similar limitations at the article level, indicating 135M models are too small for zero-shot text comprehension. Topic classification results are challenging to interpret due to small, variable data; individual dataset results should generally be interpreted cautiously. For extractive QA, results vary based on the setup. On the shuffled dataset, transfer offers limited gains. On the harder GPT-4–picked answers however, transfer improves scores, especially for 360M models.
Full LoRA 135M 360M 135M 360M En 60.43 68.13 60.43 72.52 +Da 72.52 75.82 61.53 76.92 +Is 63.73 75.82 60.43 69.23 +NoB 75.82 71.42 60.43 72.52 +NoN 76.92 74.72 67.03 71.42 +Sv 70.32 68.13 60.43 69.23 Mergeeq 75.82 72.52 60.43 62.63 Mergeis+ 68.13 80.21 60.43 63.73 Merge2 76.92 75.82 62.63 65.93
(a) Binary sentiment analysis (Sentences).
Baseline: 60.43 (135M), 60.43 ( 360M)
Full LoRA 135M 360M 135M 360M 60.71 70.23 60.71 73.80 61.90 71.42 60.71 70.23 61.90 71.42 60.71 71.42 63.09 69.04 60.71 70.23 65.47 63.09 64.28 69.04 70.23 65.47 60.71 63.09 61.90 63.09 60.71 60.71 64.28 69.04 60.71 63.09 69.04 73.80 63.09 61.90
(b) Bin. sentiment (Articles).
BL: 59.52 (135M), 60.71 ( 360M)
Full LoRA 135M 360M 135M 360M 77.35 84.18 67.94 67.94 75.21 82.47 71.79 70.94 79.05 78.20 71.79 76.06 68.80 80.76 72.22 74.78 60.68 79.05 69.65 61.11 63.67 73.50 65.38 80.34 66.23 77.77 68.37 70.94 75.64 82.90 59.82 73.93 75.21 78.63 57.26 72.64
(c) Topic Classification (Articles).
BL: 54.70 (135M), 60.25 ( 360M)
Full LoRA 135M 360M 135M 360M En 75.35 84.37 69.57 86.40 +Da 72.86 86.74 69.13 82.93 +Is 75.64 85.46 71.80 86.87 +NoB 72.04 84.16 70.00 86.60 +NoN 72.89 85.05 74.41 86.12 +Sv 70.82 85.12 70.42 86.10 Mergeeq 63.47 83.87 67.43 78.08 Mergeis+ 70.06 83.33 70.63 85.77 Merge2 72.25 85.14 70.19 82.43
(d) Extractive QA (Shuffled DS).
Baseline: 63.05 (135M), 68.94 (360M).
Full LoRA 135M 360M 135M 360M 54.40 55.25 52.40 54.85 57.65 66.70 59.70 66.35 49.20 65.05 55.35 67.35 55.85 59.15 54.35 64.60 54.65 62.60 54.30 60.45 56.20 63.45 53.00 65.25 53.20 63.70 50.85 60.05 56.60 62.05 51.85 62.70 56.35 67.65 58.75 66.00
(e) Extractive QA (LLM gen.).
BL: 32.90 (135M), 35.40 (360M).
Full LoRA 135M 360M 135M 360M 65.64 72.43 62.21 71.10 68.02 76.63 64.57 73.47 65.89 75.19 64.01 74.18 67.12 72.90 63.54 73.74 66.12 72.90 65.93 69.62 66.24 71.13 61.98 72.80 64.12 72.19 61.55 66.48 66.94 75.50 60.68 69.84 69.95 76.21 62.38 69.78
(f) Mean of the individual FoBCoMP scores in 2(a)-2(e).
Table 2: Text Comprehension Probes on datasets included in the FoBCoMP benchmark; Debess et al. (2024) and FoQA. Percentage of samples where a higher probability was assigned to the correct than to the incorrect sample.
Full LoRA 135M 360M 135M 360M En 73.61 58.27 73.61 58.27 +Da 50.98 40.45 63.23 44.38 +Is 38.77 30.09 40.05 30.54 +NoB 48.59 44.81 56.25 44.51 +NoN 61.58 50.28 68.02 56.62 +Sv 69.96 57.02 78.22 61.43 Mergeeq 103.45 42.40 182.21 338.30 Mergeis+ 94.95 40.16 68.81 62.88 Merge2 65.28 40.39 78.46 146.20
(a) Before continuing training on Faroese (zero-shot).
Full LoRA 135M 360M 135M 360M 4.98 3.75 5.51 4.48 4.19 3.56 4.25 3.55 4.44 3.48 4.53 3.53 4.22 3.63 4.26 3.56 4.60 3.66 4.90 4.08 4.26 3.60 4.21 3.58 4.08 3.41 4.61 3.93 4.08 3.41 4.58 3.77 4.22 3.49 4.56 3.80 arX
(b) After continuing training on Faroese.
Table 3: Average per-token perplexity on the Fineweb-2 evaluation set.
4.2 RQ1: Transfer Languages are Important
Across all benchmarks, initializing Faroese models with a Scandinavian transfer language improves performance compared to English-only models. The choice of transfer language, however, matters.
Perplexity
Models adapted via a Scandinavian transfer language consistently show lower perplexities than those adapted directly from English. The difference is larger for smaller models and for LoRA models, suggesting that LoRA benefits more from better-initialized parameters, consistent with Biderman et al. (2024) who find full tuning more sample-efficient than LoRA in domain adaptation.
Table 3(a) shows zero-shot perplexities before Faroese adaptation. Icelandic performs best, followed by Danish and Norwegian-Bokmål; Swedish is barely better than English-only. This likely reflects script overlap: Faroese shares most letters with Icelandic, some diacritics with Danish and Norwegian, but none beyond the English alphabet with Swedish. After Faroese adaptation (Table 3(b)), Icelandic remains best for larger models, while for smaller models, higher-resource mainland Scandinavian languages outperform Icelandic (except for Norwegian-Nynorsk, the lowest-resorce language).
FoBLiMP
While Icelandic does not always yield the best language modeling performance (perplexity after Faroese tuning), it achieves the highest scores on MultiBlimp (subject-verb agreement), as expected since Icelandic has subject-verb agreement, unlike the mainland Scandinavian languages.
Across all linguistic probes, results are mixed. Icelandic shows a small advantage (Figure 3), but it is not the top language for most task-model pairs. Mean scores across FoBLiMP tasks (Table 1(f)) show Icelandic performs best in 2 of 4 setups, making it the best individual language but not an undisputed leader. In particular, models adapted first to Icelandic do not consistently outperform those adapted to Danish or Norwegian Bokmål. Nynorsk performs worst, with no wins (Figure 3) and the lowest mean scores (Table 1(f)), likely due to limited data. This suggests that other features can compensate for lower surface similarity: the higher-resource mainland Scandinavian generally match Icelandic (except in SVA).
FoBCoMP
While Icelandic remains a strong transfer language, Nynorsk has the same win count (4), despite poor results in linguistic probes. Three of Nynorsk’s wins are from the LoRA 135M model, while other Nynorsk models perform worse, as reflected in the mean scores over FoBCoMP tasks (Table 2(f)). Table 2(f) shows Danish surpassing Icelandic in FoBCoMP: Danish wins against Icelandic in 3 of 4 aggregated cases, Nynorsk in 2/4, Bokmål and Swedish in 1/4, and English-only never. We conclude that results are very mixed.
In general,
Icelandic often provides the strongest transfer. It gives the best perplexity before Faroese adaptation and clear advantages in FoBLiMP probes like subject-verb agreement, supporting our choice of Icelandic for human evaluation. Figure 1 shows that choosing Icelandic over English gives almost consistent improvements for FoBLiMP and in many cases gains for FoBCoMP.
Figure 1: Win rates of Icelandic over English only. FoBLiMP (left), FoBCoMP (right).
Figure 2 illustrates how often all transfer setups outperform English, indicating that any additional transfer language is better than none. For FoBLiMP, especially in 360M models, transfer is crucial, while for FoBCoMP, the best language choice is less clear.
Figure 2: Win rates of all transfer setups over English only (i.e., any Scandinavian language is better than none. FoBLiMP (left), FoBCoMP (right).
Comparing single languages for the transfer (Figure 3), the choice is less clear: Icelandic has 14 wins and Danish 13. For FoBLiMP, Icelandic leads with 9 wins versus 7, and 4 ties; for FoBCoMP, Danish leads with 9 wins versus 6, and 5 ties.
Figure 3: Win rates of transfer languages across models and adaptation setups. In case of a tie, we count both. FoBLiMP (left), FoBCoMP (right).
4.3 RQ2: Effects of Merging
Perplexities
We see in Table 3(b) that merging all transfer languages (setups Mergeeq and Mergeis+) leads to the lowest perplexities in full-parameter fine-tuning, suggesting that with sufficient learning capacity — and at least for language modelling — models can benefit from a mixture of related languages. For LoRA, however, the opposite holds: the fewer languages merged, the better the results, with the setups without merging performing best. Full fine-tuning appears more robust to language merging. Merging LoRA-fine-tuned models may generally be more problematic than merging fully fine-tuned models, as other work has also shown, possibly due to a weaker degree of representation alignment (Stoica et al., 2024). Zero-shot perplexities (Table 3(a)) in particular for Mergeeq are exploded; the more is merged the higher the perplexities. But this, interestingly, still results in a better initialization for full fine-tuning.
FoBLiMP
In Table 1(f), we see that merges are beneficial for the full tuning setup: For the 135M model, the best 3 models are the 3 merges. For the 360M model, Icelandic is best, but followed by 2 merges. For LoRA however, the situation is different: For the 135M model, merges are within the same range as models with individual languages, while for the 360M models, the scores of merges are lower than for 3 out of 5 single languages.
FoBCoMP
The Merge2 setup, which combines Icelandic and Danish equally, has four wins in Table 2(f) and is clearly the best of the merges. Full fine-tuning benefits particularly the Merge2 setup, which is best overall setup for 135M and runner-up for 360M. For LoRA, single-language setups however outperform merges. It is better than Icelandic only for both full fine-tuning models, while Icelandic only is better for both LoRA models, again demonstrating that merges show promise for full fine-tuning but are less suitable for LoRAs.
Can Merging Improve Specific Aspects?
In the preceding paragraphs, we found that merges are sometimes competitive but most often outperformed by individual languages: We found that Icelandic tends to give stronger gains on the linguistic probes in FoBLiMP, while Danish seems more helpful for comprehension probes in FoBCoMP. This raises the question of whether the two languages can complement each other, with Icelandic adding linguistic strengths and Danish adding text comprehension skills. We therefore ask: Do merges add linguistic capabilities to transfer from Danish only? To test this, we compare Danish-only with merges on FoBLiMP (Table 1). Comparing each merge pairwise against Danish within otherwise equal setups, the merge wins in 30/60 cases and Danish in 29 cases, and in one case, there is a tie. To see the potential of merging, we also compare the best merge against Danish. The best merge outperforms Danish in 14/20 cases, while Danish performs better in 6 cases. These results suggest that merging can indeed add linguistic capabilities to Danish models. Exploring this further is a promising direction for future work.
Finally, we ask: Do merges add comprehension capabilities to Icelandic models? For FoBCoMP, we make a pairwise comparison of Icelandic with the merged models (based on Table 2). Overall, the merges do not perform better: Icelandic wins in 38 out of 60 comparisons, while the merges win 16 times, with 6 ties. Focusing only on the best merge in each case, the results are evenly split, with 10 wins for the merge and 10 wins for Icelandic. This suggests that merges can sometimes match or improve on Icelandic only, but the potential is smaller than for linguistic capabilities.
Human Ratings
in the sentence continuation task, (Table 5) show that the models trained on Icelandic only and the Mergeis+ model are very close: For LoRA, Mergeis+ achieves the highest overall score (3.436 versus 3.396), while for full tuning, the Icelandic model achieves higher average scores (3.413 versus 3.361). Icelandic LoRA performs best on the lexical level but is weaker in grammar, semantics and fluency. Interestingly, for semantics, the merged models score slightly higher in both cases, which is partially in line with results previously discussed in this section, where the Icelandic models were comparatively weak on comprehension tasks. This indicates that that while lexical knowledge can be effectively acquired through pre-training on Icelandic, other skills can benefit from the broader exposure provided by the merging approach. In the summarization task however (Table 5), scores of the Icelandic model are higher in both cases and across both linguistic quality and task completion, which could indicate the opposite.
Model Overall Lexical Grammar Semantics Fluency LoRA-Mergeis+ $3.436$ $3.865$ $3.645$ $3.015$ $3.216$ LoRA-Is+ $3.396$ $3.983$ $3.587$ $2.891$ $3.126$ Full-Is+ $3.413$ $3.940$ $3.600$ $2.955$ $3.160$ Full-Mergeis+ $3.361$ $3.924$ $3.535$ $2.934$ $3.051$
Table 4: Sentence Continuation scores (averages over both annotators). Scale 0–5, higher is better. Pearson’s correlation coefficient between annotators: $r=0.546$.
Model Task Completion Linguistic Quality Full-Is+ $3.01$ $4.11$ Full-Mergeis+ $2.92$ $3.82$ LoRA-Is+ $1.20$ $2.02$ LoRA-Mergeis+ $1.03$ $1.78$
Table 5: Summarization scores (average over both annotators). Scale 0–5, higher is better. Pearson correlation coefficient between annotators: $r=0.879$.
4.4 RQ3: Full Fine-Tuning versus LoRA
Overall, full fine-tuning yields higher results than LoRA across most evaluations, but particularly in merging setups (as discussed in 4.3). This advantage is especially visible in the text comprehension tasks. LoRA shows its relative strength in linguistic acceptability transfer, especially for 135M models.
Perplexity
Full fine-tuning consistently outperforms LoRA in reducing perplexity, showing that the increased learning capacity is crucial for core language modeling. The effect is even stronger when multiple transfer languages are merged, where LoRA consistently underperforms.
FoBLiMP
In Table 1(f), we see that for the smaller 135M model, LoRA outperforms full fine-tuning in 5 out of 6 cases for single languages, while for the 360M model the results are balanced (3 wins each). Looking at all individual tasks in Figure 4, LoRA again has a clear advantage for the 135M model (34 wins vs. 23 for full fine-tuning, with 3 ties), but for the 360M model both methods perform similarly, with full tuning having a slight edge. For merges, however, both the mean scores in Table 1(f) and the win rates in Figure 4 indicate that full fine-tuning is better overall. These findings highlight that LoRA performs surprisingly well for the 135M model, capturing syntactic transfer efficiently despite being more parameter-efficient. Interestingly, this contradicts our initial assumption that higher learning capacity is more important for acquiring linguistic skills, while preventing catastrophic forgetting would primarily benefit comprehension tasks.
Figure 4: Wins of LoRA vs. full FT for model twins on FoBLiMP. No merges (left), merges (right).
FoBCoMP
results for full tuning and LoRA are compared in Figure 5. Full tuning consistently outperforms LoRA: full tuning achieves 35 wins compared to 22 for LoRA (with 3 ties; see Figure 5). This may seem counterintuitive, as fine-grained language skills are not obviously critical for comprehension tasks. However, it is possible that full tuning provides a more stable adaptation process, which benefits performance on these datasets.
Figure 5: Wins of LoRA vs. full FT for model twins in FoBCoMP. No merges (left), merges (right).
Model size effects
Differences between full tuning and LoRA are mainly driven by the 135M model. For the 360M model, margins between both approaches are small, suggesting that larger models are more robust to adaptation method choice.
Human Ratings
for sentence continuation (Table 5) give mixed results: For merges, LoRA models score higher than fully tuned models (3.436 versus 3.361), while for Icelandic-only models, the fully tuned model scores higher (3.413 versus 3.396). However, for summarization (Table 5), the results are very clear: Fully tuned models score much higher. The fully tuned Icelandic model achieves the highest performance of all with a task completion score of 3.01 and linguistic quality score of 4.11, while LoRA models perform substantially worse, with LoRA-Mergeis+ scoring only 1.03 in task completion and 1.64 in linguistic quality. This strong difference suggests that full tuning provides a better, or at least more stable, surface for preserving linguistic skills during downstream task fine-tuning. Interpretations that full tuning benefits the higher-level organizational task of summarization may be possible but should be done cautiously, given the noise added by the task fine-tuning setup with very little data that we use for summarization.
4.5 Qualitative Observations
The low semantic scores across models highlight a limitation in adaptation for low-resource languages. While the models acquire competency in surface-level linguistics — producing valid Faroese vocabulary, maintaining grammatical structures, and achieving natural-sounding fluency — they struggle significantly with generating meaningful, coherent content. This reflects a fundamental challenge in language model training: while syntactic and lexical patterns can be learned from limited data through transfer from related languages, semantic understanding — which requires world knowledge — is a challenge, particularly in small models. The example in Figure 6 highlights this: although the models produce text that appears fluent, they often fail to convey coherent ideas, limiting their utility for applications requiring content generation.
Many summarization outputs exhibit mixes between languages, as in the example in Figure 7. This was not the case for sentence continuation outputs, which indicates that the low-resource tuning destroyed some of the models’ linguistic abilities.
Faroese: Translation:
Input: Børn millum fimm og 11 ár verða nú koppsett fyri koronu nú Input: Children between five and 11 years are now being vaccinated for coronavirus now
Continuation: á døgninum. Síðani frísparksmeginreglurnar vórðu settar í gildi 24. apríl hevur eingin smitta verið staðfest, síðan stigtakarin til spæliskeiið - Danmark - varð settur at hava tamarhald á korona-smittuni týsdag Continuation: during the day. Since the free kick regulations were put into effect on April 24th, no infection has been confirmed, since the initiator of the play workshop - Denmark - was set to have control over the corona infection on Tuesday
Figure 6: Example: Semantic nonsense in outputs.
Example of summary output:
Samandráttur er ein lokaliserendre i fjórðhálsparafjöllum fyrir fjölkvangna forfælja í Føroyskaflokkum til víkjandi barna uppaling er lutfalsliga sterk í Føroyum.
Figure 7: Language mixing in outputs. The model starts in Faroese, then uses language similar to Danish or Norwegian (in italics), then language resembling Icelandic (underlined), then, again, Faroese.
4.6 Limitations and Future Work
The small amount of evaluation data for Faroese is a major limitation, and even the human study covers relatively few samples. Summarization results in particular should be viewed with caution as the small validation set made hyperparameter tuning difficult, and in particular the LoRA models may not have been trained optimally.
Valuable directions for future work are to investigate how these findings change with model sizes and after instruction tuning, and if adaptive or data-driven merging strategies can lead to better transfer from multiple languages. Multilingual studies could reveal whether the same patterns hold for other low-resource languages in other families.
5 Conclusion
We studied how to adapt small LLMs to Faroese by transferring from related Scandinavian languages, merging models trained on these languages, and comparing full and parameter-efficient fine-tuning. We found that transfer from related languages is essential, but the best source varies. Icelandic gave the strongest gains in linguistic accuracy, while Danish was more helpful for comprehension. This suggests that practical applications should draw on multiple sources rather than relying only on the closest relative. Merging showed potential, although the benefits were inconsistent and more sensitive under LoRA than full fine-tuning. This indicates that merging can add complementary strengths, but requires careful design. LoRA proved effective for improving linguistic acceptability, while full fine-tuning performed better on comprehension-heavy tasks. Full fine-tuning also provided a stronger base for downstream fine-tuning on summarization in low-resource settings. Overall, our results show that transfer from related languages is key for adapting LLMs to low-resource settings, and that the optimal tuning strategy depends on priorities and target application.
Acknowledgments
This research was supported by TrustLLM funded by Horizon Europe GA 101135671. The computations were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre and by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Appendix A Training Details
Parameter Value
Optimizer AdamW (Loshchilov and Hutter, 2019)
Scheduler Cosine with 5% warmup
Batch size 256 (effective)
Context window 8192 tokens
Learning rate Full fine-tuning: $5\times 10^{-4}$
LoRA: $8\times 10^{-4}$
(also tested: $5\times 10^{-5}$$1\times 10^{-3}$)
Hardware 1 node, 4 or 8 A100 40GB GPUs
LoRA rank 256
LoRA $\alpha$ 512
LoRA #parameters 57.5M (135M); 102M (360M)
Training epochs 5 (Faroese corpus), 1 (all other languages)
Total compute 5,000 A100 (40GB) hours
Table 6: Training hyperparameters and setup for continued pre-training.
Parameter Value
Optimizer AdamW (Loshchilov and Hutter, 2019)
Scheduler Cosine with 0.1 warmup
(also tested without scheduler)
Batch size 8
Context window 8192 tokens
Learning rate $5\times 10^{-5}$
(also tested: $5\times 10^{-4}$$1\times 10^{-3}$)
Dropout 0.1
(also tested: 0)
Hardware 1 A100 40GB GPU
LoRA rank 16
(also tested: 8)
LoRA $\alpha$ 32
(also tested: 16)
Training epochs 50 (for the full dataset, 151 samples)
Tuning split 135 training / 16 validation samples
Prompting Faroese, minimalistic setup indicating start of text and summary
Table 7: Hyperparameters and setup for summarization fine-tuning experiments in the human evaluation.

### v2

@@
 Abstract
-We investigate how to adapt small, efficient LLMs to Faroese, a low-resource North Germanic language.
-Starting from English models, we continue pre-training on related Scandinavian languages, either individually or combined via merging, before fine-tuning on Faroese.
-We compare full fine-tuning with parameter-efficient tuning using LoRA, evaluating their impact on both linguistic accuracy and text comprehension.
-Due to the lack of existing Faroese evaluation data, we construct two new minimal-pair benchmarks from adapted and newly collected datasets and complement them with human evaluations by Faroese linguists.
-Our results demonstrate that transfer from related languages is crucial, though the optimal source language depends on the task: Icelandic enhances linguistic accuracy, whereas Danish boosts comprehension.
-Similarly, the choice between full fine-tuning and LoRA is task-dependent: LoRA improves linguistic acceptability and slightly increases human evaluation scores on the base model, while full fine-tuning yields stronger comprehension performance and better preserves model capabilities during downstream fine-tuning.
+We investigate strategies for adapting small, efficient language models to Faroese, a low-resource North Germanic language.
+Starting from English-pretrained models, we apply continued pre-training on related Scandinavian languages—individually or combined via model merging—before fine-tuning on Faroese.
+We compare full fine-tuning with parameter-efficient adaptation via LoRA, assessing their effects on general language modeling performance, linguistic accuracy, and text comprehension.
+To address the lack of existing Faroese evaluation resources, we construct two new minimal-pair probing benchmarks, one for linguistic acceptability and one for text comprehension, and complement them with human evaluations conducted by native Faroese linguists.
+Our results show that transfer from related languages is essential, but the optimal source language is task-dependent: Icelandic improves linguistic accuracy, while Danish boosts reading comprehension.
+The choice of adaptation method likewise depends on the target task: LoRA yields stronger linguistic acceptability and marginally higher human evaluation scores, whereas full fine-tuning produces better comprehension performance and more robust downstream fine-tuning.
+Merging multiple related languages under full fine-tuning (but not LoRA) improves general language modeling, though its benefits in the linguistic acceptability and comprehension probes are less consistent.
 1 Introduction
@@
 (2024), and even basic comprehension may fail Court and Elsner (2024).
-Coverage of these languages typically requires the largest available models, if it works at all.
-In this paper, we focus on Faroese, a North Germanic language spoken by around 70,000 people, mostly in the Faroe Islands.
+Adequate coverage of these languages typically requires the largest available models, if it works at all.
+In this paper, we focus on Faroese, a North Germanic language spoken by approximately 70,000 people, primarily in the Faroe Islands.
 Training data is scarce, with only 95 million words available in the deduplicated FineWeb-2 dataset Penedo et al.
 (2024).
-Faroese is particularly interesting for studying transfer from related languages.
-In the Nordic context, it has been described as the central Nordic language Torp (1998), reflecting its typological position in relation to the other Scandinavian languages.
-Derived from Old Norse, shaped by historical development and sociopolitical circumstances, modern Faroese shares characteristics with all the other Scandinavian languages.
-We investigate methods for adapting small generative models to Faroese through continued pre-training, with particular emphasis on leveraging transfer from related languages.
-Prior work has demonstrated that syntactic similarity is a strong predictor of transfer success Chang et al.
-(2024), and for Faroese, encoder models benefit notably from transfer from Icelandic and other Scandinavian languages Snæbjarnarson et al.
+Faroese is typologically well-suited for studying cross-lingual transfer: descended from Old Norse and shaped by a distinctive historical and sociopolitical trajectory, it has been described as the central Nordic language Torp (1998), sharing structural features with all other Scandinavian languages to varying degrees.
+We investigate methods for adapting small generative models to Faroese through continued pre-training, with particular emphasis on transfer from related languages.
+Prior work has shown that syntactic similarity is a strong predictor of transfer success Chang et al.
+(2024), and Faroese encoder models benefit notably from Icelandic and other Scandinavian languages Snæbjarnarson et al.
 (2023).
-Beyond direct transfer, we explore parameter merging (Wortsman et al., 2022; Ilharco et al., 2023; Yadav et al., 2023), which enables the combination of models fine-tuned on different languages.
-This approach allows us to control the relative influence of each source language, a flexibility that would be more costly to achieve via multilingual training.
-Our goal is to balance the close linguistic similarity of Icelandic with the larger data resources available for mainland Scandinavian languages.
-To this end, we merge models trained on various Scandinavian languages and subsequently fine-tune the merged models on Faroese.
-We also compare two fine-tuning strategies for the continued pre-training: full fine-tuning and parameter-efficient fine-tuning with LoRA (Hu et al., 2022).
-Prior evidence suggests that full fine-tuning provides higher accuracy, while LoRA better preserves previously learned skills and diversity in text generation (Biderman et al., 2024).
-Based on this, we hypothesize that full fine-tuning will yield stronger linguistic performance in Faroese, but with a higher risk of losing general reasoning and knowledge abilities.
-A major challenge for low-resource languages is the scarcity of evaluation data.
-To address this, we take two steps.
-First, we introduce two minimal-pair evaluation suites for small LLMs, combining pre-existing, newly collected, and adapted datasets: FoBLiMP, which targets syntactic and linguistic acceptability, and FoBCoMP, which targets text comprehension.
-Second, we perform expert human evaluations (carried out by Faroese linguists) on two tasks, one with and one without low-resource downstream task fine-tuning.
-Our results indicate that transfer from related languages is crucial.
-Transfer from Icelandic yields the best performance in the linguistic probes in FoBLiMP, whereas higher-resource mainland Scandinavian languages contributes more to the text comprehension probes in FoBCoMP.
-Language merging shows promise for combining these advantages, but improvements are not consistent.
-Contrary to our expectations, LoRA outperforms full fine-tuning on linguistic acceptability in both automatic and human evaluations.
-In contrast, full fine-tuning achieves better results on comprehension tasks and coupled with downstream fine-tuning.
-Our research questions are the following:
+Beyond single-language transfer, we explore parameter merging (Wortsman et al., 2022; Ilharco et al., 2023; Yadav et al., 2023) as a means of combining models trained on different languages.
+This allows fine-grained control over the relative influence of each source language—a flexibility that would be costly to achieve through multilingual training—and enables us to balance the closer morphological and lexical proximity of Icelandic against the syntactic similarity and larger data resources of the mainland Scandinavian languages.
+We also compare two adaptation strategies: full fine-tuning and parameter-efficient fine-tuning with LoRA (Hu et al., 2022).
+Prior work suggests that full fine-tuning yields higher task accuracy, while LoRA better preserves previously acquired skills and generation diversity (Biderman et al., 2024).
+Our initial hypothesis was therefore that full fine-tuning will produce stronger Faroese linguistic performance, but at a greater risk of degrading general reasoning and knowledge capabilities.
+A persistent challenge for low-resource languages is the scarcity of evaluation data.
+We address this in two ways.
+First, we introduce two minimal-pair evaluation suites: FoBLiMP, targeting linguistic acceptability, and FoBCoMP, targeting text comprehension, both combining pre-existing, newly collected, and adapted datasets.
+[^1] Both benchmarks are released on the HuggingFace Hub under identifiers jekunz/FoBLiMP and jekunz/FoBCoMP.
+Second, we conduct expert human evaluations by native Faroese linguists on two generation tasks, one with and one without downstream fine-tuning.
+To summarize, our research questions are:
   1. RQ1
-    What is the effect of transfer from (to various degrees) related languages?
+    What is the effect of transfer from languages with varying typological proximity?
   2. RQ2
-    Does merging various related languages have a benefit over choosing the closest neighbor?
+    Does merging multiple related languages offer advantages over selecting the single closest neighbor?
   3. RQ3
-    What are the differences between fine-tuning all parameters of the base model and performing adaptation with LoRA?
-2 Related Work
-Language Adaptation
-Post-hoc adaptation transfers knowledge from a high-resource or multilingual model to a new language, often using parameter-efficient methods.
+    How does full fine-tuning of all parameters compare to LoRA-based adaptation?
+Our results confirm that transfer from related languages is essential.
+Icelandic transfer yields the strongest performance on linguistic probes, while higher-resource mainland Scandinavian languages contribute more to the comprehension probes.
+Model merging shows promise as language modeling scores improve for full fine-tuning, though gains are not consistent for the more fine-grained probes.
+Contrary to our initial hypothesis, LoRA outperforms full fine-tuning on linguistic acceptability in both automatic and human evaluations; full fine-tuning performs better on comprehension tasks and when fine-tuned on text summarization.
+2 Background
+2.1 Faroese Typology
+Faroese belongs to the Insular Scandinavian branch together with Icelandic, while Norwegian, Swedish, and Danish form Mainland Scandinavian.
+This classification reflects differences in lexicon, morphology, and syntax, although Faroese also shares substantial traits with the Mainland languages Thráinsson et al.
+(2012).
+Lexically, Faroese shares cognates with both Icelandic and Danish, though no quantitative study exists Jacobsen (2021, 2022).
+Morphologically, it is closest to Icelandic, with some overlap with Norwegian Torp (1998).
+Syntactically, however, Faroese aligns more closely with Mainland Scandinavian languages Ussery and Petersen (2023); Debess (2017); Petersen (2010); Petersen and Heycock (2017); Sandøy (2005).
+Faroese therefore occupies an intermediate position: morphologically and lexically closest to Icelandic, but syntactically closer to the Mainland languages.
+This mixed profile complicates the identification of a single closest relative and makes cross-lingual transfer particularly relevant.
+To quantify Faroese’s position within the Scandinavian family, we compute lexical and syntactic similarity measures.
+Faroese is not represented in major typological databases and lacks proper vector-based representations in multilingual toolkits such as lang2vec (Littell et al., 2017; Dryer and Haspelmath, 2013), therefore we rely on alternative resources.
+Language Lexical Syntactic
+Danish 0.167 0.814
+Icelandic 0.321 0.803
+Norwegian, Bokmål 0.180 0.812
+Norwegian, Nynorsk 0.205 0.790
+Swedish 0.181 0.794
+Table 1: Lexical and Syntactic Similarity of Faroese to other Scandinavian languages.
+Lex.
+Sim.: arithmetic mean of N-gram overlap and normalized Levenshtein similarity (computed on Islex lemma pairs).
+Synt.
+Sim.: arithmetic mean of three metrics: POS-tag distribution overlap, POS-tag N-gram overlap, and POS-sequence normalized Levenshtein similarity (computed on Flores-200).
+For lexical similarity, we use the pan-Nordic dictionary Islex Úlfarsdóttir (2014).
+We compute normalized Levenshtein distance and character N-gram overlap between Faroese lemmas and their equivalents in other Scandinavian languages; the overall score is the unweighted mean of these measures; results are given in Table 1.
+While this metric does not explicitly model cognate relations or orthographic differences, it reproduces the hierarchy suggested in the linguistic literature and confirms Icelandic as lexically closest.
+For syntactic similarity, we use the parallel splits of Flores-200 Costa-jussà et al.
+(2022).
+A total of 2007 parallel sentences were POS-tagged using Stanza Qi et al.
+(2020).
+We compute three tag-based metrics: global POS distribution overlap, POS N-gram overlap, and normalized Levenshtein similarity on POS sequences; the final score is their unweighted mean.
+All languages show high syntactic similarity to Faroese, with Danish being the closest.
+The results support existing linguistic descriptions: Icelandic is morphologically and lexically closest to Faroese, while Mainland Scandinavian languages show stronger syntactic similarity.
+This typological profile motivates our experimental design, where we consider all Scandinavian languages, with particular emphasis on Icelandic.
+2.2 Language Adaptation
+Language adaptation refers to updating a pre-trained model to improve its performance in a target language through continued pre-training on target-language data.
 Early work focused on encoder models (Pfeiffer et al., 2020; Ansell et al., 2022; Ebrahimi and Kann, 2021).
-Yong et al.
-(2023) found that full fine-tuning works best for smaller models (e.g., 560M), while adapters are better for larger ones (up to 7.1B), but despite using a generative model, they do not evaluate on generative tasks.
+For generative models, Yong et al.
+(2023) found that full fine-tuning works best for smaller models (e.g., 560M parameters), while adapter-based methods are preferable for larger ones (up to 7.1B); however, despite using a generative model, they do not evaluate on generative tasks.
 Razumovskaia et al.
-(2024) show that continued pre-training with LoRA (similar to our setup) improves linguistic quality of generations, though not few-shot ability.
-Instruction tuning on translated data is another option for language adaptation (Muennighoff et al., 2023b), but it is less reliable for small languages, and even correct translations can produce “translationese” (Gellerstam, 1986).
+(2024) show that continued pre-training with LoRA improves the linguistic quality of generations, though not few-shot task performance; a setup closely related to ours.
+Kunz (2025) test different parameter-efficient methods for language adaptation, finding that LoRA placed in the feed-forward layers was the most promising; a setup which we follow in this work.
+While work on parameter-efficient language adaptation is relatively sparse, there is extensive research on parameter-efficient task adaptation and instruction fine-tuning.
+A recurring theme is the interaction between PEFT and catastrophic forgetting.
+Whitehouse et al.
+(2024) show that LoRA outperforms full fine-tuning in low-data and zero-shot cross-lingual settings, where full fine-tuning causes models to lose the ability to generate in non-training languages.
+Biderman et al.
+(2024) observe a similar pattern in domain adaptation, and Vu et al.
+(2022) find that prompt tuning is less prone to defaulting to the source language in zero-shot cross-lingual transfer.
+Beyond forgetting, Ghosh et al.
+(2024) report that full fine-tuning leads to greater memorization of the instruction-tuning dataset and thus more hallucination, whereas LoRA encourages generalization from pre-training knowledge.
+Shuttleworth et al.
+(2025) further show that LoRA forgets less even when overall performance matches full fine-tuning, indicating that the effect is not simply due to underfitting.
+However, they also find that with large learning rates, LoRA can sometimes forget more than full fine-tuning.
+It remains unclear whether these findings extend to language adaptation and how LoRA’s restricted parameter space affects adaptation success.
 Selecting effective transfer languages is crucial for language adaptation.
 Chronopoulou et al.
-(2023) show that sharing adapter parameters across related languages improves machine translation for low-resource languages.
-Similarly, Faisal and Anastasopoulos (2022) find that linguistically informed adapter designs benefit unseen languages.
-In machine translation, transfer success often depends more on source corpus size and subword overlap than on broader linguistic similarity (Lin et al., 2019).
+(2023) show that sharing adapter parameters across related languages improves translation for low-resource languages, and Faisal and Anastasopoulos (2022) find that linguistically informed adapter designs benefit unseen languages.
+In translation, syntactic similarity is predictive of transfer performance (Lee et al., 2022), although transfer success may depend more on source corpus size and subword overlap than on broader linguistic similarity (Lin et al., 2019).
 Pairing a low-resource language with a typologically similar, higher-resource language can outperform using all available languages (Neubig and Hu, 2018).
-However, they focus on translation from a low-resource language; a setup that does not require generation in the small language.
-Faroese Typology and Its Implications
-Faroese is classified as Insular Scandinavian with Icelandic, while Norwegian, Swedish, and Danish are Mainland Scandinavian.
-This division reflects differences in lexicon, morphology, and syntax, though Faroese also shares traits with Mainland Scandinavian languages Thráinsson et al.
-(2012).
-Lexically, Faroese shares words and cognates with Icelandic and Danish, but no quantitative analysis exists Jacobsen (2021, 2022).
-Morphologically, it is more similar to Icelandic, with some overlap with Norwegian Torp (1998).
-Syntactically, Faroese is closer to Mainland Scandinavian languages Ussery and Petersen (2023); Debess (2017); Petersen (2010); Petersen and Heycock (2017); Sandøy (2005).
-Because of this mixed profile, identifying a single closest neighbour language is difficult.
-In our experiments, we fine-tune using data from all Scandinavian languages, with particular attention to Icelandic, to capture the different typological influences.
-Evaluation for low-resource languages and small models
-is challenging because there is little task-specific data.
-On the other hand, few-shot evaluations often focus on knowledge-intensive tasks or tasks that require structured outputs, making them out of scope for small models.
-Perplexity is the simplest measure of a model’s fit to a language, and only requires held-out text.
-In low-resource languages, however, fine-grained tokenization from common tokenizers reduces its usefulness (Oh and Schuler, 2024).
-An alternative is information parity, which compares the negative log-likelihood of a target-language text to its English version (Tsvetkov and Kipnis, 2024).
-This metric correlates with downstream performance but requires parallel data.
-Another option is translated benchmarks, but these miss culture-specific content (Chen et al., 2024), introduce translation errors, and produce “translationese”, making the target language artificially easier for the model.
-Minimal pairs consist of two sentences that differ slightly, with one correct and one incorrect.
-Models are expected to assign higher probability to the correct sentence, capturing implicit preferences (Marvin and Linzen, 2018; He et al., 2025).
+Notably however, these studies focus on translation from a low-resource language, a setup that does not require generation in the low-resource language itself.
+2.3 Merging
+Model merging provides an efficient alternative to multi-task training by combining several fine-tuned models into a single checkpoint, often improving generalization.
+In simple linear merging (Wortsman et al., 2022; Choshen et al., 2022), the parameters of the models are averaged using scalar weights $\alpha_{i}$ that determine how much each model influences the final result.
+Task Arithmetic (Ilharco et al., 2023) instead operates on task vectors, defined as the difference between a fine-tuned model and a shared base model.
+These vectors represent the updates introduced by each task.
+The merged model is formed by adding a weighted combination of these task vectors back to the base model, again with weight coefficients $\alpha_{i}$.
+TIES (Yadav et al., 2023) further reduces interference between tasks.
+First, it introduces a density hyperparameter $p\in(0,1]$ that determines how many parameters from each task vector are retained.
+Second, TIES resolves sign conflicts across models.
+For each parameter, it determines the dominant update direction across tasks and enforces this consensus direction.
+The final merged model is obtained by adding the weighted, sparsified, and sign-aligned task vectors back to the base model.
+For task fine-tuning, merged models often preserve base model performance and have been shown to generalize better to unseen tasks; merging models fine-tuned on multiple tasks separately matches or exceeds joint multi-task training (Yadav et al., 2024).
+But as with PEFT methods, the effect of merging languages rather than tasks is less explored.
+Aakanksha et al.
+(2024) show that models safety-aligned on one language at a time and then merged outperforms data mixing.
+For continued pre-training however, Glocker et al.
+(2025) find that merged models’ performance is lower than the performance of the monolingual base models.
+A key challenge in merging LoRA models is parameter interference, where updates from different models conflict in the shared parameter space.
+Stoica et al.
+(2024) trace this to LoRAs being less aligned than fully fine-tuned models, and propose a weight alignment procedure as a remedy.
+Tang et al.
+(2024) instead propose linearizing the LoRA modules prior to merging, which they show improves post-merge performance.
+Zhao et al.
+(2025) address interference by clustering LoRAs by similarity, merging within clusters, and ensembling the results.
+Zhang and Zhou (2025) take a different approach, constraining LoRA subspaces to be orthogonal during fine-tuning so that the resulting models are more mergeable by construction.
+2.4 Low-Resource Evaluation
+Evaluation for low-resource languages and small models poses a fundamental challenge: task-specific data is scarce, and few-shot evaluations tend to target knowledge-intensive tasks poorly suited to small models.
+Perplexity offers the simplest measure of language model fit, requiring only held-out text.
+However, coarse tokenization inflates scores for low-resource languages in ways that do not reflect true linguistic competence (Oh and Schuler, 2024).
+Information parity addresses this by comparing the negative log-likelihood of a target-language text to that of its English translation (Tsvetkov and Kipnis, 2024); while this metric correlates with downstream performance, it requires parallel data.
+Translated benchmarks are another common recourse, but they miss culture-specific content (Chen et al., 2024), introduce translation artifacts, and produce translationese—unnatural target-language text that models may find spuriously easier to process.
+Minimal pairs consist of two near-identical sentences, one grammatical and one not, with models expected to assign higher probability to the correct sentence (Marvin and Linzen, 2018; He et al., 2025).
 Classical examples include subject-verb agreement (Linzen et al.
-(2016): “The key is on the table.” vs.
-“The key are on the table.”) and negative polarity items (Marvin and Linzen (2018): “No students have ever lived here.” vs.
-“Most students have ever lived here.”).
-BLiMP (Warstadt et al., 2020) provides a wide range of syntactic minimal pairs, with multilingual extensions like MultiBLiMP (Jumelet et al., 2025) including Faroese.
-COMPS (Misra et al., 2023) tests semantic knowledge (e.g., “A robin can fly.” vs.
-“A penguin/table can fly.”), and its multilingual extension (He et al., 2025) exists, though it does not cover Faroese.
-Recent work for Faroese has focused on machine translation Scalvini and Debess (2024); Scalvini et al.
-(2025a); Simonsen and Einarsson (2024); Debess et al.
-(2025), combining automatic and human assessments.
-Automatic metrics such as BLEU or chrF prove shallow and fail to capture Faroese-specific nuances Scalvini et al.
-(2025b).
-Embedding-based metrics, such as BERTScore via FoBERT Snæbjarnarson et al.
-(2023), are emerging but limited and rarely validated against human judgment.
-Human evaluation remains essential.
-Beyond MT, small Faroese evaluation datasets exist for sentiment analysis Debess et al.
-(2024) and question answering Simonsen et al.
-(2025).
+2016: “The key is | are on the table.”) and negative polarity items (Marvin and Linzen 2018: “No | Most students have ever lived here.”).
+BLiMP (Warstadt et al., 2020) provides broad syntactic coverage, with its multilingual extension MultiBLiMP (Jumelet et al., 2025) including Faroese.
+COMPS (Misra et al., 2023) tests semantic knowledge, but its multilingual extension (He et al., 2025) does not cover Faroese.
+Recent work on Faroese has focused primarily on machine translation (Scalvini and Debess, 2024; Scalvini et al., 2025a; Simonsen and Einarsson, 2024; Debess et al., 2025), where automatic metrics such as BLEU and chrF fail to capture language-specific nuances (Scalvini et al., 2025b).
+Embedding-based metrics via FoBERT (Snæbjarnarson et al., 2023) are emerging but rarely validated against human judgment, which remains essential.
+Beyond translation, small datasets exist for sentiment analysis (Debess et al., 2024) and question answering (Simonsen et al., 2025).
 3 Experimental Setup
-In this section, we introduce our experimental setup including model training (Section 3.1), merging (Section 3.2), collection of data for the automatic evaluation (Section 3.3), and the human evaluation process and data (Section 3.4).
-[^1] We make models and data available in a huggingface collection: huggingface.co/collections/jekunz/faroese-adaptation-68d6aea566b16d4f57180682.
 3.1 Training
@@
 We train for 5 epochs on the Faroese corpus (following Muennighoff et al.
-(2023a)’s scaling law for data-constraint training); we do not repeat data for other languages.
+(2023)’s scaling law for data-constraint training); we do not repeat data for other languages.
 Training details can be found in Table 7 in Appendix A.
@@
 We perform sequential continued pre-training for RQ1, first on an individual transfer language, then on the Faroese.
-We do not merge the data of the source languages with the Faroese data because this may introduce language mixing issues (Li et al., 2025).
+We do not merge the data of the source languages with the Faroese data because this may introduce language confusion and language mixing in model outputs (Li et al., 2025).
+[^2] We make models and data available in a HuggingFace collection.
 3.2 Merging
-Merging provides an efficient alternative to multi-task training by combining different fine-tuned models into one checkpoint, often improving generalization (Wortsman et al., 2022; Yadav et al., 2024).
-Task Arithmetic merging (Ilharco et al., 2023) computes task vectors as the difference between a fine-tuned model and its base, averages them, and adds the result back to the base model.
-TIES (Yadav et al., 2023) refines this by keeping only the most influential parameters, resolving sign conflicts, and using a disjoint mean to reduce interference.
-We apply TIES to languages instead of tasks, using Mergekit (Goddard et al., 2024): starting from SmolLM, we fine-tune on each language to obtain language vectors, merge them, and train the resulting model on Faroese.
-Since we continue pre-training on Faroese after merging and have limited resources, we select three promising merges that cover different language mixes and weightings: Mergeeq where we merge all five models equally (with pre-normalization weight 1 and density 0.5), Mergeis+ where we merge with bias towards Icelandic (weight 1 for Icelandic, 0.5 for all others, density 0.5), and Merge2, where we merge only two models: Icelandic and Danish (both with weight 1 and density 0.5), as Danish is the mainland Scandinavian language with the lowest perplexity after Faroese continued pre-training (see Table 3(b)).
+We apply TIES merging to the continued pre-trained models in different languages, using Mergekit (Goddard et al., 2024): starting from SmolLM, we fine-tune on each language to obtain language vectors, merge them, and train the resulting model on Faroese.
+We select three promising merges that cover different language mixes and weightings: Mergeeq where we merge all five models equally (with pre-normalization weight 1 and density 0.5), [^3] See Section 2.3 for definitions of merging parameters.
+resulting in a strong bias towards the mainland Scandinavian languages as four of them are included in the merge; Mergeis+ where we merge with bias towards Icelandic (weight 1 for Icelandic, 0.5 for all others, density 0.5), and Mergeda+is, where we merge only two models: Icelandic and Danish (both with weight 1 and density 0.5), as Danish is the mainland Scandinavian language with the lowest perplexity after Faroese continued pre-training (see Table 5).
 3.3 Automatic Evaluation
-We evaluate the perplexity on the validation set of the Faroese portion of Fineweb-2 both zero-shot and after continued pre-training on Faroese.
+We evaluate the perplexity on the validation set of the Faroese portion of Fineweb-2.
 In addition, we introduce two benchmarks: FoBLiMP for linguistic acceptability probes, and FoBCoMP for text comprehension probes.
@@
 This collection is called FoBLiMP (Faroese Benchmark of Linguistic Minimal Pairs).
+Table 9 in Appendix B provides an overview of sources, subsets and modifications, and statistics.
 To evaluate subject-verb agreement, we use the Faroese portion of MultiBLiMP (Jumelet et al., 2025), containing 232 sentences.
@@
 Concatenating all subsets gives 552 pairs for flip_neighbours and 601 pairs for delete.
-GermDetect Michael and Horbach (2025) provide automatically corrupted sentences with verb placement errors.
+GermDetect (Michael and Horbach, 2025) provide automatically corrupted sentences with verb placement errors.
 After removing pairs with no corruption, we obtain 2,026 pairs.
@@
 We introduce a set of five probes, called FoBCoMP (Faroese Benchmark of Text Comprehension Minimal Pairs).
+Table 10 in Appendix B provides an overview with additional statistics.
 We adapt the Faroese news sentiment dataset (Debess et al., 2024) (original labels: positive, negative, neutral) into minimal pairs by adding a sentiment-bearing sentence (“Hetta er gott/ringt”).
@@
 3.4 Human Evaluation
-As we do not have evaluation sets to assess fine-grained properties for Faroese generation, we perform a human evaluation to assess output quality.
-The latter requires subjective judgment of linguistic quality, naturalness, and contextual appropriateness.
-As human evaluation is time-intensive, we focus on four 360M models, which perform among the best.
-These models vary along two dimensions: full tuning versus LoRA, and transfer from Icelandic versus Mergeis+.
-The evaluators are two native Faroese speakers, both trained linguists with extensive experience in NLP model evaluation.
+Since no fine-grained evaluation sets exist for Faroese generation, we conduct a human evaluation to assess output quality across dimensions requiring subjective judgment: linguistic quality, naturalness, and contextual appropriateness.
+Given the resource-intensive nature of expert evaluation, we focus on four 360M models representing the best-performing configurations, varying along two dimensions: full fine-tuning versus LoRA, and transfer from Icelandic versus Mergeis+.
+Two native Faroese speakers serve as evaluators, both trained linguists with extensive NLP evaluation experience.
+We first followed a structured calibration protocol: evaluators jointly examined sample outputs, discussed edge cases, and iteratively refined the scoring guidelines until reaching consensus.
+Following this calibration phase, evaluators independently annotated a small set of approximately 10 samples, then met to resolve disagreements and clarify ambiguities before proceeding independently.
+Periodic check-ins were held throughout to maintain consistency without sharing individual scores.
+Task Dimension $r$ $\kappa_{w}$ 95% CI Sent.
+Cont.
+Sem.
+coh.
+0.427 0.383 [0.303, 0.460] Lex.
+cor.
+0.409 0.286 [0.212, 0.357] Grammar 0.375 0.367 [0.277, 0.449] Fluency 0.298 0.295 [0.205, 0.379] Summ.
+Task compl.
+0.728 0.718 [0.644, 0.781] Ling.
+qual.
+0.829 0.819 [0.760, 0.866]
+Table 2: Inter-annotator agreement: Pearson’s $r$ and quadratic-weighted Cohen’s $\kappa_{w}$ with bootstrapped 95% confidence intervals (10,00010{,}000 iterations).
+Sent.
+cont.: $n=400$, summ.: $n=200$.
 Sentence Continuation
-Given that we are working with small base models, we chose a simple generation task: sentence continuation.
-Models are prompted with sentences from a small manually compiled corpus derived from academic papers and local news (not in FineWeb2), with the last words removed and a trailing space.
-All models produce running text, enabling comparison of output quality.
-We evaluate outputs for linguistic quality across four subdimensions, scored 0–5: Lexical correctness (valid Faroese words and avoidance of lexical hallucinations), grammatical accuracy (morphological and syntactic correctness, including spelling and typography), Semantic coherence (meaningful, logically consistent content) and fluency/naturalness (native-like expression).
-The annotators each evaluated 400 continuations (100 prompts, 4 models) We apply a token cut-off of 100 and no penalty for incomplete last sentences.
+Given the limited capacity of small base models, we adopt sentence continuation as a straightforward generation task.
+Models are prompted with sentences drawn from a small manually compiled corpus of academic papers and local news articles (not included in FineWeb-2), with the final words removed and a trailing space added.
+All models produce running text, enabling direct comparison of output quality.
+A token cut-off of 100 was applied.
+Outputs are evaluated along four subdimensions of linguistic quality, each scored 0–5: lexical correctness (valid Faroese vocabulary, absence of hallucinated forms), grammatical accuracy (morphological and syntactic well-formedness, including spelling and typography), semantic coherence (meaningful, internally consistent content), and fluency/naturalness (native-like expression).
+Each annotator evaluated 400 continuations (100 prompts $\times$ 4 models).
+During guideline development, annotators agreed to rate the full generated output rather than the immediate completion alone, as the truncated sentence was typically resolved within a few words before the model continued generating further text.
+Inter-annotator agreement, measured by Pearson’s $r$, was 0.546 overall; all scores are reported in Table 2.
 Summarization
-As zero-shot summarization did not result in reasonable summaries, we create a synthetic dataset for fine-tuning and manually evaluate the summaries generated by the models.
-The synthetic dataset is compiled with 150 authentic texts in the same domains as the evaluation set (academic, news, blog) paired with summaries generated by Claude Sonnet 4222For fine-tuning details, see Table 7 in Appendix A..
-For evaluation, we selected 50 source texts (not part of training text) and generated summaries from all four models, resulting in 200 summary-source pairs.
-Two annotators conducted a blind evaluation using two criteria: task completion (to what extent did the model solve the task) and linguistic quality, each on a 0–5 scale.
-We apply a token cut-off of 400 and no penalty for incomplete final sentences.
+Initial zero-shot experiments (i.e., prompting models directly to summarize a given text) failed to produce recognizable summaries, necessitating fine-tuning.
+As there was no Faroese summarization dataset, we constructed a synthetic one: 150 authentic texts spanning the same domains as the evaluation set (academic, news, blog), each paired with a summary generated by Claude Sonnet 4.444Fine-tuning details are provided in Table 8 in Appendix A.
+For evaluation, we selected 50 source texts not in FineWeb-2 and generated summaries from all four models (token cutoff: 400), yielding 200 summary–source pairs.
+Annotators rated two criteria: task completion (the extent to which the model successfully performed the summarization task) and linguistic quality, each rated on a 0–5 scale.
+Guidelines were developed through the same iterative calibration process as for sentence continuation, though the criteria were intentionally broader, focusing on overall task success and linguistic quality rather than fine-grained subdimensions.
+One point of disagreement during calibration concerned the threshold for assigning a score of 0, which required additional discussion to resolve.
+Inter-annotator agreement was substantially higher than for sentence continuation (Pearson’s r=0.879 overall; Table 2), likely because the coarser-grained criteria reduced room for subjective interpretation.
 4 Results and Discussion
@@
 Models perform well on most linguistic acceptability probes, suggesting these tasks are relatively easy.
-The main exception is Translation Pairs, where scores are lower, likely due to noise as translation error counts do not always reflect linguistic quality: an error can reflect unrelated aspects such as incorrect content compared to the source sentence.
-Full LoRA 135M 360M 135M 360M En 95.25 96.55 96.12 96.55 +Da 96.98 95.25 98.27 96.12 +Is 97.41 99.13 97.41 98.27 +NoB 96.12 100.0 96.98 96.98 +NoN 96.12 98.27 95.68 97.41 +Sv 96.12 98.27 96.98 96.98 Mergeeq 97.41 98.27 96.55 95.25 Mergeis+ 97.84 97.84 97.41 96.55 Merge2 97.84 97.41 97.41 99.13
+The main exception is Translation Pairs, where scores are lower, possibly as translation error counts do not always reflect linguistic quality: translation errors can reflect aspects such as incorrect content compared to the source sentence.
+Full LoRA 135M 360M 135M 360M En 95.25 96.55 96.12 96.55 +Da 96.98 95.25 98.27 96.12 +Is 97.41 99.13 97.41 98.27 +NoB 96.12 100.0 96.98 96.98 +NoN 96.12 98.27 95.68 97.41 +Sv 96.12 98.27 96.98 96.98 Mergeeq 97.41 98.27 96.55 95.25 Mergeis+ 97.84 97.84 97.41 96.55 Mergeda+is 97.84 97.41 97.41 99.13
 (a) Subject-verb agreement (MultiBLiMP).
@@
 BL: 65.39 (135M), 67.38 (360M)
-Full LoRA 135M 360M 135M 360M En 93.48 95.36 92.54 94.66 +Da 94.27 95.11 94.91 95.75 +Is 94.27 95.16 94.57 95.75 +NoB 95.06 95.31 94.91 95.55 +NoN 93.97 94.91 94.47 95.01 +Sv 94.61 95.85 95.36 95.80 Mergeeq 94.66 95.80 95.16 95.31 Mergeis+ 94.57 95.60 95.06 95.26 Merge2 94.52 95.85 94.37 95.60
+Full LoRA 135M 360M 135M 360M En 93.48 95.36 92.54 94.66 +Da 94.27 95.11 94.91 95.75 +Is 94.27 95.16 94.57 95.75 +NoB 95.06 95.31 94.91 95.55 +NoN 93.97 94.91 94.47 95.01 +Sv 94.61 95.85 95.36 95.80 Mergeeq 94.66 95.80 95.16 95.31 Mergeis+ 94.57 95.60 95.06 95.26 Mergeda+is 94.52 95.85 94.37 95.60
 (d) Verb placement (GermDetect).
@@
 Full LoRA 135M 360M 135M 360M 88.00 90.92 86.78 89.65 90.48 91.12 90.86 91.49 89.65 92.25 90.39 92.80 90.20 91.78 90.65 91.98 88.59 91.34 88.87 90.46 90.43 91.55 90.87 90.99 90.56 91.73 90.55 90.89 91.10 91.98 90.22 90.81 90.57 91.81 90.28 91.45
-(f) Mean of the individual FoBLiMP scores in 1(a)-1(e).
+(f) Mean of the individual FoBLiMP scores in 3(a)-3(e).
 Table 3: Linguistic probes on datasets included in FoBLiMP: Percentage of samples where a higher probability was assigned to the original than to the corrupted sample.
@@
 On the harder GPT-4–picked answers however, transfer improves scores, especially for 360M models.
-Full LoRA 135M 360M 135M 360M En 60.43 68.13 60.43 72.52 +Da 72.52 75.82 61.53 76.92 +Is 63.73 75.82 60.43 69.23 +NoB 75.82 71.42 60.43 72.52 +NoN 76.92 74.72 67.03 71.42 +Sv 70.32 68.13 60.43 69.23 Mergeeq 75.82 72.52 60.43 62.63 Mergeis+ 68.13 80.21 60.43 63.73 Merge2 76.92 75.82 62.63 65.93
+Full LoRA 135M 360M 135M 360M En 60.43 68.13 60.43 72.52 +Da 72.52 75.82 61.53 76.92 +Is 63.73 75.82 60.43 69.23 +NoB 75.82 71.42 60.43 72.52 +NoN 76.92 74.72 67.03 71.42 +Sv 70.32 68.13 60.43 69.23 Mergeeq 75.82 72.52 60.43 62.63 Mergeis+ 68.13 80.21 60.43 63.73 Mergeda+is 76.92 75.82 62.63 65.93
 (a) Binary sentiment analysis (Sentences).
@@
 BL: 54.70 (135M), 60.25 ( 360M)
-Full LoRA 135M 360M 135M 360M En 75.35 84.37 69.57 86.40 +Da 72.86 86.74 69.13 82.93 +Is 75.64 85.46 71.80 86.87 +NoB 72.04 84.16 70.00 86.60 +NoN 72.89 85.05 74.41 86.12 +Sv 70.82 85.12 70.42 86.10 Mergeeq 63.47 83.87 67.43 78.08 Mergeis+ 70.06 83.33 70.63 85.77 Merge2 72.25 85.14 70.19 82.43
+Full LoRA 135M 360M 135M 360M En 75.35 84.37 69.57 86.40 +Da 72.86 86.74 69.13 82.93 +Is 75.64 85.46 71.80 86.87 +NoB 72.04 84.16 70.00 86.60 +NoN 72.89 85.05 74.41 86.12 +Sv 70.82 85.12 70.42 86.10 Mergeeq 63.47 83.87 67.43 78.08 Mergeis+ 70.06 83.33 70.63 85.77 Mergeda+is 72.25 85.14 70.19 82.43
 (d) Extractive QA (Shuffled DS).
@@
 Full LoRA 135M 360M 135M 360M 65.64 72.43 62.21 71.10 68.02 76.63 64.57 73.47 65.89 75.19 64.01 74.18 67.12 72.90 63.54 73.74 66.12 72.90 65.93 69.62 66.24 71.13 61.98 72.80 64.12 72.19 61.55 66.48 66.94 75.50 60.68 69.84 69.95 76.21 62.38 69.78
-(f) Mean of the individual FoBCoMP scores in 2(a)-2(e).
+(f) Mean of the individual FoBCoMP scores in 4(a)-4(e).
 Table 4: Text Comprehension Probes on datasets included in the FoBCoMP benchmark; Debess et al.
@@
 Percentage of samples where a higher probability was assigned to the correct than to the incorrect sample.
-Full LoRA 135M 360M 135M 360M En 73.61 58.27 73.61 58.27 +Da 50.98 40.45 63.23 44.38 +Is 38.77 30.09 40.05 30.54 +NoB 48.59 44.81 56.25 44.51 +NoN 61.58 50.28 68.02 56.62 +Sv 69.96 57.02 78.22 61.43 Mergeeq 103.45 42.40 182.21 338.30 Mergeis+ 94.95 40.16 68.81 62.88 Merge2 65.28 40.39 78.46 146.20
-(a) Before continuing training on Faroese (zero-shot).
-Full LoRA 135M 360M 135M 360M 4.98 3.75 5.51 4.48 4.19 3.56 4.25 3.55 4.44 3.48 4.53 3.53 4.22 3.63 4.26 3.56 4.60 3.66 4.90 4.08 4.26 3.60 4.21 3.58 4.08 3.41 4.61 3.93 4.08 3.41 4.58 3.77 4.22 3.49 4.56 3.80 arX
-(b) After continuing training on Faroese.
-Table 3: Average per-token perplexity on the Fineweb-2 evaluation set.
+Full LoRA 135M 360M 135M 360M En 4.98 3.75 5.51 4.48 +Da 4.19 3.56 4.25 3.55 +Is 4.44 3.48 4.53 3.53 +NoB 4.22 3.63 4.26 3.56 +NoN 4.60 3.66 4.90 4.08 +Sv 4.26 3.60 4.21 3.58 Mergeeq 4.08 3.41 4.61 3.93 Mergeis+ 4.08 3.41 4.58 3.77 Mergeda+is 4.22 3.49 4.56 3.80
+Table 5: Average per-token perplexity on the Fineweb-2 evaluation set after CPT on Faroese.
 4.2 RQ1: Transfer Languages are Important
+Figure 1: Win rates of transfer languages across models and setups. In case of a tie, we count both.
 Across all benchmarks, initializing Faroese models with a Scandinavian transfer language improves performance compared to English-only models. The choice of transfer language, however, matters.
@@
 (2024) who find full tuning more sample-efficient than LoRA in domain adaptation.
-Table 3(a) shows zero-shot perplexities before Faroese adaptation.
-Icelandic performs best, followed by Danish and Norwegian-Bokmål; Swedish is barely better than English-only.
-This likely reflects script overlap: Faroese shares most letters with Icelandic, some diacritics with Danish and Norwegian, but none beyond the English alphabet with Swedish.
-After Faroese adaptation (Table 3(b)), Icelandic remains best for larger models, while for smaller models, higher-resource mainland Scandinavian languages outperform Icelandic (except for Norwegian-Nynorsk, the lowest-resorce language).
+Table 5) shows that after adaptation on Faroese, Icelandic is the best individual transfer language for larger models, while for smaller models, higher-resource Mainland Scandinavian languages outperform Icelandic (except for Norwegian-Nynorsk, the lowest-resource Mainland Scandinavian language).
 FoBLiMP
@@
 We conclude that results are very mixed.
-In general,
-Icelandic often provides the strongest transfer.
-It gives the best perplexity before Faroese adaptation and clear advantages in FoBLiMP probes like subject-verb agreement, supporting our choice of Icelandic for human evaluation.
-Figure 1 shows that choosing Icelandic over English gives almost consistent improvements for FoBLiMP and in many cases gains for FoBCoMP.
-Figure 1: Win rates of Icelandic over English only. FoBLiMP (left), FoBCoMP (right).
-Figure 2 illustrates how often all transfer setups outperform English, indicating that any additional transfer language is better than none.
-For FoBLiMP, especially in 360M models, transfer is crucial, while for FoBCoMP, the best language choice is less clear.
-Figure 2: Win rates of all transfer setups over English only (i.e., any Scandinavian language is better than none. FoBLiMP (left), FoBCoMP (right).
-Comparing single languages for the transfer (Figure 3), the choice is less clear: Icelandic has 14 wins and Danish 13.
-For FoBLiMP, Icelandic leads with 9 wins versus 7, and 4 ties; for FoBCoMP, Danish leads with 9 wins versus 6, and 5 ties.
-Figure 3: Win rates of transfer languages across models and adaptation setups. In case of a tie, we count both. FoBLiMP (left), FoBCoMP (right).
+Discussion
+Icelandic emerges as the strongest transfer language overall.
+It yields the lowest perplexity prior to Faroese adaptation and shows clear advantages on FoBLiMP probes such as subject–verb agreement, supporting our decision to use Icelandic for human evaluation.
+Compared to English-only initialization, Icelandic provides almost consistent improvements on FoBLiMP (19/20 cases) and frequent gains on FoBCoMP (13/20 cases).
+We next examine how often all transfer setups outperform English, which would indicate that any additional transfer language is preferable to none.
+For FoBLiMP, transfer is particularly important for the 360M models: all transfer languages outperform English in 9/10 cases for 360M models, compared to 4/10 cases for 135M models.
+For FoBCoMP, however, the picture is less clear, with all transfer setups outperforming English in only 4/20 cases.
+When comparing all individual transfer languages (compare Figure 1), Icelandic achieves 14 wins overall and Danish 13.
+On FoBLiMP, Icelandic leads (9 wins vs.
+7; 4 ties), whereas on FoBCoMP Danish performs better (9 wins vs.
+6; 5 ties).
+The overall finding that initializing Faroese models with a Scandinavian transfer language improves performance compared to English-only models confirms prior results for Faroese encoder models (Snæbjarnarson et al., 2023) and extends them to decoder models.
+Snæbjarnarson et al.
+(2023) report that an Icelandic BERT base model often outperforms a Danish base model, consistent with our results for perplexity and linguistic acceptability.
+Interestingly, they find that Danish BERT performs better on a semantic text similarity benchmark (the only semantic task in their work), aligning with our observation that Danish shows a slight tendency to outperform Icelandic on text comprehension tasks.
+More broadly, our findings are consistent with prior work on language transfer, which highlights both the importance of transfer languages and the benefits of using typologically related languages (Chronopoulou et al., 2023; Faisal and Anastasopoulos, 2022; Chang et al., 2024).
 4.3 RQ2: Effects of Merging
@@
 Merging LoRA-fine-tuned models may generally be more problematic than merging fully fine-tuned models, as other work has also shown, possibly due to a weaker degree of representation alignment (Stoica et al., 2024).
-Zero-shot perplexities (Table 3(a)) in particular for Mergeeq are exploded; the more is merged the higher the perplexities.
+Zero-shot perplexities before adaptation on Faroese (Table 11(a); Appendix C) are extremely high for merges, in particular for Mergeeq; the more is merged, the higher the perplexities.
 But this, interestingly, still results in a better initialization for full fine-tuning.
@@
 For the 360M model, Icelandic is best, but followed by 2 merges.
-For LoRA however, the situation is different: For the 135M model, merges are within the same range as models with individual languages, while for the 360M models, the scores of merges are lower than for 3 out of 5 single languages.
+For LoRA however, the situation is different: For the 135M model, merges are within the same range as models with individual languages, while for the 360M models, the scores of merges are lower than for 3 out of 5 individual languages.
 FoBCoMP
-The Merge2 setup, which combines Icelandic and Danish equally, has four wins in Table 2(f) and is clearly the best of the merges.
-Full fine-tuning benefits particularly the Merge2 setup, which is best overall setup for 135M and runner-up for 360M.
+The Mergeda+is setup, which combines Icelandic and Danish equally, has four wins in Table 4(f) and is clearly the best of the merges.
+Full fine-tuning benefits particularly the Mergeda+is setup, which is best overall setup for 135M and runner-up for 360M.
 For LoRA, single-language setups however outperform merges.
@@
 Can Merging Improve Specific Aspects?
-In the preceding paragraphs, we found that merges are sometimes competitive but most often outperformed by individual languages: We found that Icelandic tends to give stronger gains on the linguistic probes in FoBLiMP, while Danish seems more helpful for comprehension probes in FoBCoMP.
-This raises the question of whether the two languages can complement each other, with Icelandic adding linguistic strengths and Danish adding text comprehension skills.
-We therefore ask: Do merges add linguistic capabilities to transfer from Danish only?
-To test this, we compare Danish-only with merges on FoBLiMP (Table 1).
-Comparing each merge pairwise against Danish within otherwise equal setups, the merge wins in 30/60 cases and Danish in 29 cases, and in one case, there is a tie.
-To see the potential of merging, we also compare the best merge against Danish.
-The best merge outperforms Danish in 14/20 cases, while Danish performs better in 6 cases.
-These results suggest that merging can indeed add linguistic capabilities to Danish models.
-Exploring this further is a promising direction for future work.
-Finally, we ask: Do merges add comprehension capabilities to Icelandic models?
-For FoBCoMP, we make a pairwise comparison of Icelandic with the merged models (based on Table 2).
-Overall, the merges do not perform better: Icelandic wins in 38 out of 60 comparisons, while the merges win 16 times, with 6 ties.
-Focusing only on the best merge in each case, the results are evenly split, with 10 wins for the merge and 10 wins for Icelandic.
-This suggests that merges can sometimes match or improve on Icelandic only, but the potential is smaller than for linguistic capabilities.
+Our preceding analyses indicate that merges are sometimes competitive but often outperformed by individual languages.
+Icelandic tends to drive stronger gains on linguistic probes (FoBLiMP), whereas Danish appears more beneficial for comprehension probes (FoBCoMP).
+This suggests that merges could potentially leverage the linguistic strengths of Icelandic while incorporating the comprehension skills of Danish.
+We find that merges can indeed add linguistic capabilities to Danish models.
+Comparing Danish-only models with merges on FoBLiMP (Table 3), pairwise comparisons show that merges outperform Danish in 30 out of 60 cases, Danish wins in 29 cases, and one case ties.
+Considering the best merge against Danish, the merge achieves 14 wins out of 20, compared to 6 for Danish.
+These results indicate that merging can meaningfully enhance the linguistic capabilities of Danish models.
+For comprehension capabilities, merges provide only limited benefits to Icelandic models.
+On FoBCoMP (Table 4), in 60 pairwise comparisons, Icelandic wins 38 times, merges 16, with 6 ties.
+Focusing only on the best merge per case, results are evenly split (10 wins each).
+This suggests that while merges can occasionally match or slightly improve comprehension over Icelandic-only models, the gains are smaller than those observed for linguistic capabilities.
 Human Ratings
@@
 In the summarization task however (Table 6), scores of the Icelandic model are higher in both cases and across both linguistic quality and task completion, which could indicate the opposite.
-Model Overall Lexical Grammar Semantics Fluency LoRA-Mergeis+ $3.436$ $3.865$ $3.645$ $3.015$ $3.216$ LoRA-Is+ $3.396$ $3.983$ $3.587$ $2.891$ $3.126$ Full-Is+ $3.413$ $3.940$ $3.600$ $2.955$ $3.160$ Full-Mergeis+ $3.361$ $3.924$ $3.535$ $2.934$ $3.051$
-Table 4: Sentence Continuation scores (averages over both annotators). Scale 0–5, higher is better. Pearson’s correlation coefficient between annotators: $r=0.546$.
-Model Task Completion Linguistic Quality Full-Is+ $3.01$ $4.11$ Full-Mergeis+ $2.92$ $3.82$ LoRA-Is+ $1.20$ $2.02$ LoRA-Mergeis+ $1.03$ $1.78$
-Table 5: Summarization scores (average over both annotators). Scale 0–5, higher is better. Pearson correlation coefficient between annotators: $r=0.879$.
+Overall Lex. Gram. Sem. Fluency LoRA-Mergeis+ 3.436 3.865 3.645 3.015 3.216 LoRA-Is 3.396 3.983 3.587 2.891 3.126 Full-Is 3.413 3.940 3.600 2.955 3.160 Full-Mergeis+ 3.361 3.924 3.535 2.934 3.051
+(a) Sentence Continuation. Pearson $r=0.546$.
+Task Compl. Ling. Quality 3.01 4.11 2.92 3.82 1.20 2.02 1.03 1.78
+(b) Summarization. $r=0.879$.
+Table 6: Results for the human evaluation. Scores are averages over annotators. Scale 0–5, higher is better.
+Discussion
+Our results confirm the observation from prior work that LoRAs are more challenging to merge than fully fine-tuned models (Stoica et al., 2024; Tang et al., 2024; Zhang and Zhou, 2025), so this outcome is consistent with expectations.
+More notably, we show that merging models fine-tuned on related languages can meaningfully benefit CPT in a new target language.
+Unlike task merging, where performance on the original tasks is generally preserved or improved (Yadav et al., 2024), previous attempts at language merging have shown that out-of-the-box performance often deteriorates (Glocker et al., 2025).
+[^5] Our out-of-the-box perplexities on Faroese FineWeb confirm this trend, with merged models performing worse than individual source languages; see Table 11(a) in Appendix C.
+In contrast, by continuing pre-training on the target language after merging, we observe gains especially in perplexity, demonstrating that while it may not lead to strong performance out-of-the-box, language merging is a very promising initialization strategy for further fine-tuning.
+We speculate that the merges lead to parameter interference, which the continued training helps resolve.
+Finally, although prior work indicates that larger models are generally easier to merge (Dang et al., 2024; Yadav et al., 2024), we do not observe a strong size effect in our CPT settings.
+This however does not preclude effects for substantially larger models, where merging may become even more impactful.
+Overall, our findings show that initializing with merges can leverage complementary capabilities across related languages, at least in the context of full fine-tuning.
 4.4 RQ3: Full Fine-Tuning versus LoRA
-Overall, full fine-tuning yields higher results than LoRA across most evaluations, but particularly in merging setups (as discussed in 4.3).
-This advantage is especially visible in the text comprehension tasks.
-LoRA shows its relative strength in linguistic acceptability transfer, especially for 135M models.
 Perplexity
@@
 FoBLiMP
-In Table 1(f), we see that for the smaller 135M model, LoRA outperforms full fine-tuning in 5 out of 6 cases for single languages, while for the 360M model the results are balanced (3 wins each).
-Looking at all individual tasks in Figure 4, LoRA again has a clear advantage for the 135M model (34 wins vs.
-23 for full fine-tuning, with 3 ties), but for the 360M model both methods perform similarly, with full tuning having a slight edge.
-For merges, however, both the mean scores in Table 1(f) and the win rates in Figure 4 indicate that full fine-tuning is better overall.
-These findings highlight that LoRA performs surprisingly well for the 135M model, capturing syntactic transfer efficiently despite being more parameter-efficient.
-Interestingly, this contradicts our initial assumption that higher learning capacity is more important for acquiring linguistic skills, while preventing catastrophic forgetting would primarily benefit comprehension tasks.
-Figure 4: Wins of LoRA vs. full FT for model twins on FoBLiMP. No merges (left), merges (right).
+In the mean results in Table 3(f), we observe that for the smaller 135M model, LoRA outperforms full fine-tuning in 5 out of 6 single-language settings.
+For the 360M model, however, performance is evenly split, with three wins each.
+Considering win rates across all individual tasks yields a similar pattern.
+For the 135M model, LoRA shows a clear advantage (22 wins vs.
+7 for full fine-tuning, with 3 ties).
+In contrast, for the 360M model the two methods perform comparably, with a slight edge for full fine-tuning (12 wins for LoRA vs.
+16 for full fine-tuning, with 2 ties).
+For merged transfer setups, both the mean scores in Table 3(f) and the overall win rates indicate an advantage for full fine-tuning.
+Across these conditions, LoRA achieves 8 wins, compared to 22 for full fine-tuning.
+These findings highlight that LoRA performs surprisingly well for individial languages in the 135M model.
+It captures syntactic transfer efficiently despite its parameter-efficiency.
+This contradicts our initial assumption that higher learning capacity is more important for acquiring linguistic skills, while preventing catastrophic forgetting would primarily benefit comprehension tasks.
 FoBCoMP
-results for full tuning and LoRA are compared in Figure 5.
-Full tuning consistently outperforms LoRA: full tuning achieves 35 wins compared to 22 for LoRA (with 3 ties; see Figure 5).
-This may seem counterintuitive, as fine-grained language skills are not obviously critical for comprehension tasks.
-However, it is possible that full tuning provides a more stable adaptation process, which benefits performance on these datasets.
-Figure 5: Wins of LoRA vs. full FT for model twins in FoBCoMP. No merges (left), merges (right).
-Model size effects
-Differences between full tuning and LoRA are mainly driven by the 135M model.
-For the 360M model, margins between both approaches are small, suggesting that larger models are more robust to adaptation method choice.
+Overall, full fine-tuning consistently outperforms LoRA, achieving 35 wins compared to 22 for LoRA (with 3 ties).
+At first glance, this may seem somewhat unexpected, as fine-grained linguistic adaptation is not obviously more crucial for text comprehension tasks than for acceptability judgments.
+However, full fine-tuning may enable a more thorough and stable adaptation of the model, which in turn benefits performance on these datasets.
+The overall difference is primarily driven by the 135M model.
+Here, full fine-tuning clearly dominates (22 wins vs.
+6 for LoRA, with 2 ties).
+For the 360M model, by contrast, the margin is small (16 wins for LoRA vs.
+13 for full fine-tuning, with 1 tie), suggesting that the larger models are less sensitive to the choice of adaptation method and generally more robust across setups.
 Human Ratings
@@
 Interpretations that full tuning benefits the higher-level organizational task of summarization may be possible but should be done cautiously, given the noise added by the task fine-tuning setup with very little data that we use for summarization.
+Discussion
+Our findings reveal a clear task-dependent pattern.
+Full fine-tuning is overall the more stable and robust adaptation method.
+Its advantage is most evident in perplexity reduction, text comprehension, and especially summarization.
+It achieves higher results than LoRA across most evaluations, but particularly in merging setups (as discussed in 4.3).
+LoRA shows its strength in linguistic acceptability, indicating that parameter-efficient adaptation is sufficient — and sometimes advantageous — for capturing syntactic regularities.
+This contrasts with our initial assumption that higher learning capacity would be most important for fine-grained and language-dependent linguistic skills: In prior work, LoRA is often associated with improved generalization and reduced catastrophic forgetting (Biderman et al., 2024; Ghosh et al., 2024).
+From this perspective, one might expect LoRA to be more beneficial for comprehension tasks, where preserving broader and more language-independent knowledge should matter most.
+Instead, we observe that full fine-tuning dominates in comprehension and summarization, while LoRA’s advantage is in linguistic acceptability.
+Previous studies report that LoRA is particularly effective and stable in low-data regimes (Whitehouse et al., 2024; Vu et al., 2022).
+Although Faroese is a low-resource language, our CPT uses a 95M-word corpus, which does not clearly fall into the low-data scenarios typically examined in previous literature that looked at task adaptation, where low-data often means samples in the hundreds.
+Nevertheless, the gains for LoRA on acceptability tasks suggest that parameter-efficient adaptation helps to better learn core syntactic regularities and to transfer them to related languages.
 4.5 Qualitative Observations
-The low semantic scores across models highlight a limitation in adaptation for low-resource languages.
+The low semantic scores across models confirm a limitation in small models, especially for low-resource languages like Faroese.
 While the models acquire competency in surface-level linguistics — producing valid Faroese vocabulary, maintaining grammatical structures, and achieving natural-sounding fluency — they struggle significantly with generating meaningful, coherent content.
-This reflects a fundamental challenge in language model training: while syntactic and lexical patterns can be learned from limited data through transfer from related languages, semantic understanding — which requires world knowledge — is a challenge, particularly in small models.
-The example in Figure 6 highlights this: although the models produce text that appears fluent, they often fail to convey coherent ideas, limiting their utility for applications requiring content generation.
+This reflects a fundamental challenge in language model training: while syntactic and lexical patterns can be learned from limited data through transfer from related languages, semantic understanding, which requires world knowledge, is a challenge.
 Many summarization outputs exhibit mixes between languages, as in the example in Figure 2.
-This was not the case for sentence continuation outputs, which indicates that the low-resource tuning destroyed some of the models’ linguistic abilities.
-Faroese: Translation:
-Input: Børn millum fimm og 11 ár verða nú koppsett fyri koronu nú Input: Children between five and 11 years are now being vaccinated for coronavirus now
-Continuation: á døgninum.
-Síðani frísparksmeginreglurnar vórðu settar í gildi 24.
-apríl hevur eingin smitta verið staðfest, síðan stigtakarin til spæliskeiið - Danmark - varð settur at hava tamarhald á korona-smittuni týsdag Continuation: during the day.
-Since the free kick regulations were put into effect on April 24th, no infection has been confirmed, since the initiator of the play workshop - Denmark - was set to have control over the corona infection on Tuesday
-Figure 6: Example: Semantic nonsense in outputs.
+This was not the case for sentence continuation outputs, which indicates that the low-resource tuning interferes with the models’ linguistic abilities.
 Example of summary output:
@@
 The model starts in Faroese, then uses language similar to Danish or Norwegian (in italics), then language resembling Icelandic (underlined), then, again, Faroese.
-4.6 Limitations and Future Work
-The small amount of evaluation data for Faroese is a major limitation, and even the human study covers relatively few samples.
-Summarization results in particular should be viewed with caution as the small validation set made hyperparameter tuning difficult, and in particular the LoRA models may not have been trained optimally.
-Valuable directions for future work are to investigate how these findings change with model sizes and after instruction tuning, and if adaptive or data-driven merging strategies can lead to better transfer from multiple languages.
-Multilingual studies could reveal whether the same patterns hold for other low-resource languages in other families.
+4.6 Generality of Results.
+To assess generalizability, we replicate our approach for Northern Sámi (Uralic) and Upper Sorbian (West Slavic), both even lower-resourced than Faroese and less related to English, the language of the base model.
+The core findings hold: typologically more related languages (Finnish/Estonian for Northern Sámi; Czech/Polish for Upper Sorbian) consistently outperform majority/contact languages (Swedish/Norwegian; German) on perplexity, confirming our finding from Section 4.2 about typological similarity as an important factor for transfer effectiveness.
+As in Section 4.3, merged models under full fine-tuning again achieve the lowest perplexities, confirming merging as a particularly successful initialization strategy for further continued pre-training on the target language.
+On linguistic acceptability, LoRA scores are even here consistently better than full fine-tuning scores, confirming the result we found surprising, that LoRA is particularly good for linguistic acceptability, from Section 4.4.
+One notable divergence to the Faroese results is that LoRA outperforms full fine-tuning on perplexity for these two languages, possibly reflecting stronger regularization benefits when adapting to smaller or noisier corpora.
+Comprehension task results remain largely inconclusive across both languages.
+Details on the setup and full results are reported in Appendix D.
 5 Conclusion
-We studied how to adapt small LLMs to Faroese by transferring from related Scandinavian languages, merging models trained on these languages, and comparing full and parameter-efficient fine-tuning.
-We found that transfer from related languages is essential, but the best source varies.
-Icelandic gave the strongest gains in linguistic accuracy, while Danish was more helpful for comprehension.
-This suggests that practical applications should draw on multiple sources rather than relying only on the closest relative.
-Merging showed potential, although the benefits were inconsistent and more sensitive under LoRA than full fine-tuning.
-This indicates that merging can add complementary strengths, but requires careful design.
-LoRA proved effective for improving linguistic acceptability, while full fine-tuning performed better on comprehension-heavy tasks.
-Full fine-tuning also provided a stronger base for downstream fine-tuning on summarization in low-resource settings.
-Overall, our results show that transfer from related languages is key for adapting LLMs to low-resource settings, and that the optimal tuning strategy depends on priorities and target application.
+This article examined strategies for adapting small LLMs to Faroese, comparing transfer from related Scandinavian languages, model merging across multiple source languages, and full versus parameter-efficient LoRA fine-tuning.
+Our results establish that transfer from related languages is essential, but no single source language dominates: Icelandic proved most valuable for linguistic accuracy, while Danish contributed more to comprehension.
+This complementarity suggests that one should draw on multiple source languages simultaneously, which motivates experiments on model merging.
+Under full fine-tuning, merging consistently improved perplexity, showing that merged models are a successful initialization strategy for continued pre-training on Faroese.
+Gains were however less reliable for finer-grained probes, and merged LoRA adapters showed lower stability.
+The choice of adaptation strategy entails a trade-off: LoRA was more effective for linguistic acceptability judgments, whereas full fine-tuning produced stronger comprehension performance and a more robust foundation for downstream tasks.
+Future work should examine if our results transfer to larger and instruction-tuned models, and explore whether adaptive or data-driven merging strategies can yield more consistent transfer gains.
+Limitations
+The scarcity of Faroese evaluation data is a fundamental limitation, both for this study and for low-resource NLP more broadly.
+Even our human evaluation covers relatively few samples, as linguist-administered rating is labor-intensive and costly.
+Summarization results in particular should be interpreted with caution: the small validation set was a much less reliable signal for hyperparameter tuning than a validation set of more significant size would be, and in particular the LoRA models, which are more sensitive to hyperparameters, may consequently not have been trained optimally.
+We experiment with only two model sizes, one LoRA configuration, and restrict ourselves to English-only base models, chosen as the cleanest experimental baseline given the absence of small (<<1B) multilingual models with well-documented training data composition.
+Extending this work to multilingual base models that include (some of) the target languages would be a valuable next step.
+The English-centric tokenizer of our base models is another limitation: it increases inference cost for non-English languages and may also affect model performance.
+For real-world deployment and to optimize performance, transferring to a multilingual tokenizer, or one specifically adapted to the Northern Germanic languages, would be preferable.
 Acknowledgments
@@
 Appendix A Training Details
+This section provides additional details to support reproducibility, supplementing Section 3.1 and Section 3.4.
+Hyperparameters for continued pre-training are reported in Table 7; those for summarization fine-tuning are reported in Table 8.
 Parameter Value
@@
 LoRA $\alpha$ 512
+LoRA modules Feed-forward layers (following Kunz (2025))
 LoRA #parameters 57.5M (135M); 102M (360M)
@@
 Table 8: Hyperparameters and setup for summarization fine-tuning experiments in the human evaluation.
+Appendix B Dataset Statistics
+This section provides tabular overviews of our two new benchmark suites introduced in Section 3.3, detailing source datasets, subsets, modifications, and example counts.
+FoBLiMP is described in Table 9; FoBCoMP in Table 10.
+Both benchmarks are publicly released on HuggingFace under identifiers jekunz/FoBLiMP and jekunz/FoBCoMP, licensed under Creative Commons licenses: FoBLiMP under Creative Commons Attribution ShareAlike 4.0, FoBCoMP under Creative Commons Attribution Non-Commercial ShareAlike 4.0; consistent with the source datasets.
+Corruption Type Source Subset Pairs Corruption Method
+– Existing Work – – – Subject-Verb Agreement MultiBLiMP (Jumelet et al., 2025) Full 232 Automatic using morphological annotation
+Verb Placement GermDetect (Michael and Horbach, 2025) All corrupted samples 2,026 Automatic using dependency trees
+Flipped Words ScaLA (Nielsen, 2023) flip_neighbors 552 Automatic using simple heuristics
+Deleted Words ScaLA (Nielsen, 2023) delete 601 Automatic using simple heuristics
+Ours – Translation Pairs New Full 680 (from 100 source sentences) Machine-translated sentences ranked after human ratings
+Table 9: Datasets and corruption methods used for the linguistic acceptability probes included in FoBLiMP.
+Task Source Subset Pairs MinPair Generation
+Sentiment Classification FoSent (Debess et al., 2024) Sentence; both annotators agree 91 Add a sentence at the end: This is good / bad.
+Sentiment Classification FoSent (Debess et al., 2024) Document; both annotators agree 84 Add a sentence at the end: This is good / bad.
+Topic Classification FoSent (Debess et al., 2024) Deemed relevant by human annotator 234 Add a sentence at the end: The topic of this article is: X.
+Exchange related topics (e.g., Local NewsInternational News).
+Extractive QA FoQA (Simonsen et al., 2025) Full 21,867 (2,000 unique texts & correct answers) Dataset shuffling (negative answer passage is also part of the context, length controlled)
+Extractive QA FoQA (Simonsen et al., 2025) Full 2,000 GPT-4-generated: extract a different, wrong answer with the same length.
+Table 10: Datasets and corruption methods used for the text comprehension probes included in FoBCoMP.
+Appendix C Extended Perplexity Results
+Full LoRA 135M 360M 135M 360M En 73.61 58.27 73.61 58.27 +Da 50.98 40.45 63.23 44.38 +Is 38.77 30.09 40.05 30.54 +NoB 48.59 44.81 56.25 44.51 +NoN 61.58 50.28 68.02 56.62 +Sv 69.96 57.02 78.22 61.43 Mergeeq 103.45 42.40 182.21 338.30 Mergeis+ 94.95 40.16 68.81 62.88 Mergeda+is 65.28 40.39 78.46 146.20
+(a) Before continuing training on Faroese.
+Full LoRA 135M 360M 135M 360M 4.98 3.75 5.51 4.48 4.19 3.56 4.25 3.55 4.44 3.48 4.53 3.53 4.22 3.63 4.26 3.56 4.60 3.66 4.90 4.08 4.26 3.60 4.21 3.58 4.08 3.41 4.61 3.93 4.08 3.41 4.58 3.77 4.22 3.49 4.56 3.80
+(b) After continuing training on Faroese.
+Table 11: Average per-token perplexity on the Fineweb-2 evaluation set, extending results in Table 5.
+See Table 11(a).
+Appendix D Generality of the Results: Experiments on Northern Sámi and Upper Sorbian
+To assess which of our Faroese findings generalize to other languages, we replicate the approach for two additional low-resource languages: Northern Sámi and Upper Sorbian.
+Both have fewer speakers and smaller corpora than Faroese and are classified by UNESCO as endangered (Northern Sámi) and vulnerable (Upper Sorbian) (Moseley, 2010).
+Neither functions as the majority language in its region, and public life is dominated by majority languages (Dagsvold et al., 2015; McMonagle, 2022; Mitschke, 2025).
+Typologically, Upper Sorbian is Indo-European (West Slavic), while Northern Sámi belongs to the Uralic family.
+This allows us to test whether the Faroese findings generalize both within Indo-European and across language families.
+[^6] Unlike the Faroese–Icelandic case, the closest transfer languages for Northern Sámi (Finnish) and Upper Sorbian (Czech and Polish) are not themselves low-resource, so we cannot test whether a more distant but higher-resource language improves semantic performance.
+D.1 Experimental Setup
+We focus on the 360M parameter models and compare full fine-tuning and LoRA using the same hyperparameters as in the Faroese experiments.
+For each target language and adaptation method, we create two merges: one combining two languages (a typologically close language and the regional majority language), and one including one or two additional related languages.
+For Northern Sámi, we use Finnish and Estonian as the related languages (Kahn and Valijärvi, 2017), and Swedish and Norwegian (Bokmål) as majority, high-contact languages.
+Most Northern Sámi speakers are bilingual in one of these languages (Kahn and Valijärvi, 2017), which influences local varieties (Jokinen et al., 2016) and results in lexical borrowing (Pietikäinen, 2008, 2010).
+For Upper Sorbian, we select Czech and Polish as West Slavic neighbors and German as the regional majority language; all Sorbs speak German, often as their dominant language (Mitschke, 2025).
+Evaluation relies on massively multilingual benchmarks due to the lack of suitable curated resources.
+We report perplexity on the respective FineWeb-2 test sets and evaluate linguistic acceptability with MultiBLiMP (Jumelet et al., 2025) (2536 pairs for Northern Sámi, 186 for Upper Sorbian) and reading comprehension with MultiWikiQA (Smart, 2025) (4146 examples for Northern Sámi, 4906 for Upper Sorbian).
+For MultiWikiQA, we follow the non-LLM negative sampling procedure used in the Faroese experiments.
+D.2 Results and Conclusions
+Table 12.
+Perplexity $\downarrow$ MultiBLiMP $\uparrow$ MultiWikiQA $\uparrow$ Transfer Full LoRA Full LoRA Full LoRA Se En 6.5180 6.7996 0.8052 0.8111 0.7786 0.7742 +Fi 6.0851 5.1539 0.8222 0.8722 0.7735 0.7726 +Ee 6.2498 5.4192 0.8261 0.8600 0.7646 0.7820 +NoB 6.2706 5.2190 0.7973 0.8324 0.7742 0.7735 +Sv 6.3907 5.4526 0.7993 0.8360 0.7670 0.7800 +Fi+Sv 6.0597 5.5669 0.8285 0.8391 0.7752 0.7656 +All 5.8965 5.8592 0.8206 0.8324 0.7742 0.7677 Hsb En 7.3967 8.5332 0.7366 0.7527 0.7790 0.7729 +Cs 6.4403 5.8660 0.8011 0.8280 0.7831 0.7927 +De 6.9620 6.5548 0.7258 0.7903 0.7650 0.7725 +Pl 6.5469 5.8850 0.7527 0.8441 0.7811 0.7915 +De+Pl 6.4882 6.5572 0.7527 0.7957 0.7644 0.7707 +All 6.3261 6.3270 0.7581 0.8280 0.7748 0.7729
+Table 12: Results for Northern Sámi and Upper Sorbian across CPT and merging configurations.
+Effect of transfer languages.
+A first continued pre-training step on a related language consistently improves perplexity compared to models without a transfer language.
+In both languages, typologically related transfer languages yield the largest gains: Finnish and Estonian outperform Norwegian Bokmål and Swedish for Northern Sámi, and Czech and Polish outperform German for Upper Sorbian.
+This replicates our Faroese finding that typological similarity predicts transfer effectiveness.
+Majority and contact languages also help, but generally less so.
+For linguistic acceptability probes (MultiBLiMP), typologically closer languages consistently improve over the English-only baseline.
+For more distant contact languages, LoRA models show consistent improvements over English while fully fine-tuned models do not, suggesting that stronger regularization helps when transfer languages are less related.
+Results on comprehension probes are less clear.
+For Northern Sámi, MultiWikiQA differences are marginal across setups.
+Upper Sorbian shows somewhat clearer improvements from related-language transfer, though gains remain small overall.
+Overall, these results confirm our Faroese findings from Section 4.2: transfer languages help, typological similarity matters, and morphological similarity is particularly important for linguistic acceptability.
+Effect of merging.
+Under full fine-tuning, merged models again achieve the lowest perplexities, confirming our finding from Section 4.3 that combining multiple transfer sources improves language modeling.
+Merged LoRA adapters, however, underperform the best single-language variants.
+For linguistic acceptability, results are mixed.
+For Northern Sámi MultiBLiMP under full fine-tuning, the Finnish–Swedish merge scores highest, while for Upper Sorbian a single language (Czech) wins.
+Under LoRA, a single language always scores highest (Finnish and Polish, respectively).
+For MultiWikiQA, merges and individual languages perform similarly for Northern Sámi under full fine-tuning, while merged LoRA adapters score lowest and Polish is the consistent winner for Upper Sorbian.
+We conclude that merging reliably helps language modeling and can help linguistic acceptability, but does not help comprehension.
+This is consistent with the Faroese results in Section 4.3.
+LoRA versus full fine-tuning.
+Unlike in the Faroese experiments (Section 4.4), where full fine-tuning slightly outperformed LoRA on perplexity, LoRA often achieves lower perplexity for Northern Sámi and Upper Sorbian.
+The gap is even larger for linguistic acceptability, where LoRA consistently outperforms full fine-tuning, especially for Upper Sorbian.
+This may reflect stronger regularization when adapting to smaller or noisier corpora.
+QA results remain inconclusive, particularly for Northern Sámi where differences across setups are minimal.